### v1

# Enhancing Sequential Recommendations through Multi-Perspective Reflections and Iteration

Abstract.
Sequence recommendation (SeqRec) is a typical task aimed at predicting the next interaction item for users, focusing on understanding user interactive intents and collaborative filtering information. Currently, large language models (LLMs) show great potential in recommendation scenarios through prompt-based, fixed reflection library, and fine-tuning approaches. However, these approaches face challenges such as lack of supervision, inability to optimize reflection sources, inflexibility to diverse users, and high computational costs.
Despite achieving encouraging performance, current studies solely exploit reflections on users’ explicit preferences (e.g., item titles). They ignore reflecting on implicit preferences and collaborative filtering information, which prevents them from capturing the preference shifts embedded in these signals, neglecting reflections on other aspects of preferences(such as implicit preferences (e.g., brands) and collaborative filtering information. As a result, it struggles to capture users’ dynamic preferences. Moreover, current methods lack reflection evaluation and iteration, which inevitably affects the quality of reflections and leads to sub-optimal recommendation. To address these issues, we propose a Mixture of REflectors framework, namely MoRE, for modeling and learning the dynamic user preferences in sequential recommendations. Specifically, we introduce three reflectors for each user to generate LLMs’ reflections on explicit user preferences, implicit user preferences, and collaborative signals. Within each reflector, we devise a self-improving strategy called refining-and-iteration to evaluate and update the reflections. On top of these reflectors, we introduce a meta-reflector responsible for selecting the most appropriate expert and the corresponding reflections for each user’s current recommendation by applying a contextual bandit algorithm. This ensures that we can capture dynamic user preferences effectively.
Extensive experiments on three real-world datasets demonstrate that our method consistently outperforms state-of-the-art approaches. Additionally, we verify that MoRE requires the least training time and GPU memory compared to all other training LLM methods in SeqRec.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06
1. Introduction
Sequential recommendation (SeqRec), which predicts the next item of interest based on a user’s historical interaction sequence, is crucial for recommender systems. The key to this task is to accurately capture users’ dynamic preferences.
Large language models (LLMs) are emerging as promising recommenders (LLMREC) due to their vast world knowledge and excellent reasoning abilities. Two typical approaches in this context are prompt-based methods and fine-tuning methods. Prompt-based methods (Dai et al., 2023; Hou et al., 2023) pre-construct fixed prompts and exploit in-context learning (Dong et al., 2022) to guide the LLM in reasoning the desired item. Fine-tuning methods (Zheng et al., 2024; Zhang et al., 2024) inject domain knowledge by fine-tuning an LLM on substantial annotated recommendation data. Although these methods achieve encouraging performance, the former suffers from an inability to optimize the prompts based on user feedback, while the latter requires substantial computational resources.
Recently, reflection-based methods (Yao et al., 2023; Wang et al., 2024) have emerged as an alternative approach. As shown in Fig. 1, these methods allow LLMs to reflect on their own outputs based on user history and in turn improve future recommendations. Despite their effectiveness, they suffer from the following challenges in modeling and learning user preferences. First, these methods typically rely on explicit preferences, such as titles, to represent items and expect that LLMs can capture user interest patterns based on these features. However, user interests are often implicitly embedded in the attributes of items. For instance, consider a user who is an avid fan of Apple products. If she purchases an Apple smartphone and an Apple t-shirt, the connection between these items is based on the brand rather than the item titles. The implicit preference, in this case, is the brand “Apple,” which signifies a deeper user preference that is not captured solely by examining the titles of the items. Additionally, collaborative filtering (CF) information, which analyzes user behaviors and preferences to find patterns and similarities between users, is not effectively utilized in current approaches. This omission leads to a loss of valuable insights that can be gained from user interactions and preferences. Second, users’ interests, hidden in their behaviors, are intrinsically dynamic and evolve over time. For example, the recommendation at a given time step may depend on reasoning based on explicit preferences, whereas the recommendation at the next step may need to focus on implicit preferences. Therefore, it is necessary to dynamically model and select explicit preferences, implicit preferences, and CF information to ensure that the recommender system can adapt to the changing nature of user preferences and provide more accurate and relevant recommendations. Third, an ideal reflection update strategy should balance both effectiveness and efficiency, but current methods only manage to achieve one of these qualities. One approach (Wang et al., 2024) pre-establishes a fixed pool of reflection candidates to choose from, which may lack the adaptability required for changing user preferences. Another approach (Yao et al., 2023) updates the reflection candidate set using large amounts of data. While this method is potentially more effective in capturing dynamic interests, it is resource-intensive and less efficient.
Figure 1. Workflow of reflection-based methods for SeqRec. LLM reflections are generated offline based on the user’s interaction history ①. These reflections are then collected ② and used as prompts for future online recommendations ③.
Facing the above challenges, we propose a Mixture of REflectors framework, namely MoRE, for modeling and learning the dynamic user preferences in SeqRec. Specifically, we introduce three reflectors for each user to collect LLMs’ reflections based on user interaction history, each focusing on different aspects: explicit user preferences, implicit user preferences, and collaborative signals. To dynamically update the generated reflections, we devise a meta-reflector with a self-improving strategy called refining and iteration. In the refining step, we evaluate the quality of the reflections by measuring the relative improvement in LLMREC, comparing scenarios with and without reflection prompts. When the improvement exceeds a certain threshold, the reflection is considered effective. In the iteration step, we feed the selected effective reflections into LLMs to generate improved reflections, which can then be sent back to the refining step for a new round of evaluation. As such, this offline, self-improving property ensures the effectiveness and efficiency of the reflection update. Moreover, the meta-reflector operates on top of three reflectors, responsible for selecting the most appropriate reflection for each user’s current recommendation. The selection process is formulated as a contextual bandit problem (Li et al., 2010), which can be efficiently optimized. Experimental results demonstrate that MoRE consistently outperforms traditional recommendation methods and LLMREC methods including prompt-based, fine-tuning-based, and reflection-based with minimal GPU memory usage and training time overhead. In summary, our contributions are as follows:
  • We propose an effective and efficient framework, MoRE, for modeling and learning dynamic user preferences in LLM-based SeqRec.
  • MoRE incorporates three reflectors, each collecting reflections from explicit user preferences, implicit user preferences, and CF signals. It also integrates a meta-reflector, which updates the reflections with a refining-and-iteration strategy and selects the appropriate reflection for LLMREC using a contextual bandit algorithm.
  • Extensive experiments on three real-world datasets demonstrate that MoRE outperforms state-of-the-art approaches in terms of recommendation performance with minimal GPU memory usage and training time overhead.
2. Related Works
2.1. Sequential Recommendation
SeqRec aims to predict a user’s next item of interest by modeling their past interactions chronologically. The core challenge lies in accurately capturing users’ dynamic preferences. FPMC (Rendle et al., 2010) is an early attempt that addressed this by combining matrix factorization with a Markov chain to capture user preferences. In recent years, deep learning methods have significantly advanced recommendation performance. For instance, Hidasi et al. (2015); Tang and Wang (2018); Li et al. (2017) employ convolutional and recurrent neural networks to model local and global interaction patterns. Meanwhile, Li et al. (2022, 2023); Zhou et al. (2022) develop sequential recommendation models based solely on multi-layer perceptrons, demonstrating comparable performance. Additionally, Transformer-based models proposed by Kang and McAuley (2018); Sun et al. (2019); He et al. (2021, 2022) offer precise item-to-item relevance modeling through self-attention, achieving state-of-the-art results.
Despite their success, these methods often struggle to fully capture the complexity and fluidity of user preferences. This limitation is due to the lack of world knowledge and comprehensive reasoning ability. Consequently, methods based on large language models have emerged as a promising option for sequential recommendation.
2.2. Large Language Models for Recommendation
Current approaches to adapting LLMs for SeqRec predominantly fall into two categories: prompt-based methods and fine-tuning methods. Prompt-based methods involve manually constructing fixed prompts to guide LLMs in reasoning toward the desired items. For example, LLM4RS (Dai et al., 2023) enhances LLM’s recommendation capabilities by customizing prompts to align with traditional ranking, while LLMRank (Hou et al., 2024) leverages LLMs for zero-shot ranking in recommendations through specialized prompting strategies. However, these fixed prompts rely heavily on human expertise and struggle to adapt to diverse users, often resulting in sub-optimal performance (Dong et al., 2022; Wang et al., 2024). Fine-tuning methods (Geng et al., 2022), on the other hand, formulate SeqRec as a question answering task and fine-tune accordingly. To incorporate recommendation knowledge, they either add special tokens (e.g., Llara (Liao et al., 2024) and LC-Rec (Zheng et al., 2024)) or insert embeddings as texts into the prompt (e.g., BinLLM (Zhang et al., 2024)). Despite these efforts, a significant gap remains between LLM’s pre-training and recommendation-oriented fine-tuning, which necessitates a substantial amount of tuning data to achieve proper alignment.
Recently, Wang et al. (2024) propose Re2LLM, a reflection-based approach that leverages LLMs to reflect on interaction history to enhance future recommendations. However, this approach maintains a static prompt pool and solely reflects on explicit user preferences.
3. MoRE: The Proposed Framework
3.1. Problem Formulation
In SeqRec, let $\mathcal{V}=\{v_{1},v_{2},...,v_{|\mathcal{V}|}\}$ denote the item set, $\mathcal{U}=\{u_{1},u_{2},...,u_{|\mathcal{U}|}\}$ denote the user set, and $\mathcal{S}_{u}=[v_{1}^{u},v_{2}^{u},...,v_{t}^{u}]$ denote the interaction sequence in chronological order for user $u\in\mathcal{U}$, where $v_{t}^{u}\in\mathcal{V}$ is the item that $u$ has interacted with at time step $t$. The goal of this task is to predict the item that $u$ will interact with at $t+1$ given $\mathcal{S}_{u}$:
(1) $\hat{v}=\operatorname*{arg\,max}_{v\in\mathcal{V}}\Pr(v_{t+1}^{u}=v\mid\mathcal{S}_{u}).$
In this study, we focus on the LLM-based recommendation. Following (Dai et al., 2023; Hou et al., 2024), we adopt LLM to directly make the prediction as follows:
(2) $\hat{\mathcal{O}}_{u}=\mathrm{LLM}_{\mathrm{REC}}(P(u,\mathcal{S}_{u},\mathcal{C}_{u})),\hat{\mathcal{O}}_{u}\subseteq\mathcal{C}_{u},$
where $P$ denotes the manually crafted recommendation prompt; $\mathcal{C}_{u}$ and $\mathcal{O}_{u}$ respectively denote the candidate and predicted item list for $u$; all the parameters are concatenated in the text form; and $\mathrm{LLM}_{\mathrm{REC}}$, the LLM for recommendation, is frozen. Please note that reflections are essentially a form of prompts generated by the LLM based on interaction history. Therefore, replacing $P()$ with $Ref()$ in Eq. 2 denotes the reflection-based recommendation.
3.2. Overall Framework
We propose MoRE, a novel reflection-based framework, to model the dynamic preferences in LLM-based SeqRec. As shown in Fig. 2, MoRE consist of two key modules:
Figure 2. The overall architecture of our MoRE framework. MoRE incorporates multi-perspective reflectors to generate reflections offline ① and a meta-reflector to self-improve these reflections ② and learn to select the most appropriate one ③. During online recommendation ④, the meta-reflector selects the suitable reflections for the current user.
Multi-Perspective Reflectors. As discussed in Sec. 1, user preferences are often implicitly embedded in the attributes of items and can be inferred through collaborative filtering information. Therefore, MoRE incorporates this module to drive LLMs to generate reflections from the perspectives of a user’s explicit preferences, implicit preferences, and collaborative filtering information. These derived reflections based on historical interactions are considered as hints to enhance LLMs’ future recommendations. The details of this module will be introduced in Sec. 3.3.
Meta-Reflector. This module has two main objectives. First, it maintains three memory banks for each user to collect reflections from the corresponding three reflectors and uses a refining-and-iteration strategy to self-improve the reflections and update the current reflection memories for each user. Second, it is responsible for deciding which of the three perspectives of reflection should be chosen to enhance the current recommendation at each step for the user. This decision-making process is formulated as a contextual bandit problem (Li et al., 2010). These two parts will be elaborated in Sec. 3.4 and Sec. 3.5, respectively.
3.3. Multi-Perspective Reflections Generation
As illustrated in the left of Fig. 2, MoRE aims at exploiting users’ historical interaction to generate reflections from the perspectives of explicit preference, implicit preference, and collaborative signals. We achieve this goal through the following reflectors:
Explicit Preference Reflector infers user $u$’s explicit preferences (e.g., item titles) by driving a LLM to reflect the interacted items $\mathcal{S}_{u}$, the candidate items $\mathcal{C}_{u}$, and the difference between LLM’s prediction $\hat{\mathcal{O}}_{u}$ and ground-truth $\mathcal{O}_{u}$. LLM generate these reflections simultaneously and return the overall reflection $Ref_{\mathrm{EP}}$, which can be formulated as:
(3) $Ref_{\mathrm{EP}}=\mathrm{LLM}_{\mathrm{REF}}(P_{\mathrm{EP}}(u,\mathcal{S}_{u},\mathcal{C}_{u},\hat{\mathcal{O}}_{u},\mathcal{O}_{u})),$
where $\mathrm{LLM}_{\mathrm{REF}}$ is a frozen LLM specific for reflection generation, and $P_{\mathrm{EP}}$ denotes the prompt for reflecting from the perspective of explicit user preferences, which represents items using their titles. The details are shown in App. B.3.
Implicit Preference Reflector is designed to capture the preference embedded in the other item attributes (e.g., item brands) from the interaction sequences $\mathcal{S}_{u}$, the candidate set $\mathcal{C}_{u}$ and the LLM prediction $\hat{\mathcal{O}}_{u}$ and label list $\mathcal{O}_{u}$. This reflector functions similarly to the explicit preference reflector and adopts a frozen LLM to generate reflections:
(4) $Ref_{\mathrm{IP}}=\mathrm{LLM}_{\mathrm{REF}}(P_{\mathrm{IP}}(u,\mathcal{S}_{u},\mathcal{C}_{u},\hat{\mathcal{O}}_{u},\mathcal{O}_{u})),$
where $P_{\mathrm{IP}}$ denotes the reflection prompt from the perspective of implicit user preferences, representing items using their brand names, as detailed in App. B.4.
Collaborative Filtering Reflector targets guiding LLM to reflect from the perspectives of collaborative filtering information. To incorporate the CF signals, instead of introducing new token embeddings into the vocabulary (Zheng et al., 2024) or inserting special sequences into prompts (Zhang et al., 2024), we adopt a pre-trained CF model to rate the CF scores $\mathcal{M}_{u}^{\mathcal{S}}$ and $\mathcal{M}_{u}^{\mathcal{C}}$ respectively for the interacted items for $\mathcal{S}_{u}$ and the candidate items $\mathcal{C}_{u}$. These scores are then involved in the input of the reflector, which is formulated as:
(5) $Ref_{\mathrm{CF}}=\mathrm{LLM}_{\mathrm{REF}}(P_{\mathrm{CF}}(u,\mathcal{S}_{u},\mathcal{C}_{u},\mathcal{M}_{u}^{\mathcal{S}},\mathcal{M}_{u}^{\mathcal{C}},\hat{\mathcal{O}}_{u},\mathcal{O}_{u})),$
where $P_{\mathrm{CF}}$ denotes the reflection prompt from the perspective of CF, as detailed in App. B.5.
We maintain three memory banks for each user to store the generated reflections from the corresponding three perspectives. These memory banks are guided by the meta-reflector to improve the reflections, as will be elaborated in the following section.
3.4. Reflection Memory Maintenance
Considering that LLMs may not always accurately generate reflection results, we need a mechanism to ensure the validity of reflections after the meta-reflector obtains them from the three reflectors. Therefore, we define a measure called the reflection improvement effect to assess the reflections and devise a refining-and-iteration strategy to update the reflections generated by the LLM.
Reflection Improvement Effect is defined as the difference in LLM recommendation performance on the validation set, with or without the reflections. In other words, if incorporating reflections improves the LLM’s performance, it indicates that the reflection is effective for recommendation. The greater the performance improvement, the more effective the reflection. If the reflection improvement effect of a reflection exceeds a preset threshold, we regard it as an effective one. Formally, we formulate the reflection improvement effect as:
(6) $Imp$ $=\mathrm{Metric}(\hat{\mathcal{O}}_{u}^{\mathrm{REF}})-\mathrm{Metric}(\hat{\mathcal{O}}_{u}^{\mathrm{REC}})$
$\hat{\mathcal{O}}_{u}^{\mathrm{REF}}$ $=\mathrm{LLM}_{\mathrm{REC}}(Ref_{i}(u,\mathcal{S}_{u},\mathcal{C}_{u})),$
$\hat{\mathcal{O}}_{u}^{\mathrm{REC}}$ $=\mathrm{LLM}_{\mathrm{REC}}(P(u,\mathcal{S}_{u},\mathcal{C}_{u})),$
where $Ref_{i}\in\{Ref_{\mathrm{EP}},Ref_{\mathrm{IP}},Ref_{\mathrm{CF}}\}$, $\mathrm{Metric}$ denotes the recommendation metric, e.g., NDCG@10.
Refining. As we filter out reflections that offer less or negative SeqRec improvement using the above measure, we further refine the selected reflections with three strategies:
(1) Global Level: We select the reflections that bring the greatest average improvement for all users. In this approach, all users share the same reflections.
(2) Group Level: We use a collaborative filtering (CF) model to construct embeddings for all users and then cluster them using the K-means++ algorithm (Bahmani et al., 2012). We then choose the reflections that provide the greatest average improvement for users within the same cluster, ensuring that users in the same cluster share the same reflections.
(3) Individual Level: We select the reflections that bring the greatest improvement for each user. This strategy leads to personalized reflections.
All three of the aforementioned strategies are greedy, which may lead to local optima. To mitigate this issue, we propose using the improvement effect brought about by each reflection as the sampling probability. Reflections sampled from the filtered set serve as demonstrations and are exploited in the subsequent iteration step.
Iteration. Since both reflection generation and refining are based on users’ historical interactions, a major advantage of MoRE is that these two processes can be completed offline. As such, we can further use the sampled reflections as demonstrations and task the three reflectors with generating improved reflections in an in-context learning fashion (Dong et al., 2022). In this reflection generation and iteration loop, we enable the LLM to self-improve the reflections. We will further validate the impact of reflection iteration on LLM recommendation performance in the experiments.
3.5. Reflection Selection
The second objective of the meta-reflector is to model preference shifts in SeqRec. Since we have designed three memory banks for each user to collect effective reflections from three different perspectives, we require the meta-reflector to select one of these reflections for recommendation to achieve this goal. We formulate the decision-making process as a Multi-Armed Contextual Bandit (MACB) problem (Li et al., 2010), and adopt the Proximal Policy Optimization (PPO) algorithm (Schulman et al., 2017) to efficiently learn to select the reflections. Formally, we define MACB as a tuple $(\mathcal{Z},\mathcal{A},\mathcal{R})$, where $\mathcal{Z},\mathcal{A},\mathcal{R}$ denote the state space, action (arm) space, and reward function, respectively.
  • State $\bm{z}\in\mathcal{Z}$ is defined as the recommendation context. We represent it by concatenating the embedding of user $u$ and the interaction sequence $\mathcal{S}_{u}$, i.e., $\bm{z}=[emb_{u},emb_{v}]$, where the embeddings are derived from a CF model, and $emb_{v}$ denotes the average embedding for $\mathcal{S}_{u}$.
  • Action $\bm{a}\in\mathcal{A}$ is defined as the selection of a reflection from three memory banks. We represent $\bm{a}$ as a 3-dimensional vector, with each dimension corresponding to an arm, each arm representing a reflection memory.
  • Reward is defined to assess the recommendation improvement brought by the reflections. We reuse the measure in Eq.(6) and represent the reward as $R(\bm{z,a})=Imp$. Since the reflection has been assessed in Sec. 3.3 we do not need to call the LLM here. Additionally, due to the sampling discussed in Sec. 3.4, $R(\bm{z,a})$ can be both positive and negative, ensuring training efficiency.
  • Replay Buffer is designed to facilitate efficiency in policy optimization. We represent it as $D=(\bm{z},\bm{a},R(\bm{z},\bm{a}),\bm{z}^{\prime})$ to store the tuples of observed state, action, reward, and next state. With the records in the replay buffer, we can refine successful policies and learn from erroneous trials.
Training
We apply the PPO algorithm (Schulman et al., 2017) with an Actor & Critic networks and a Clip objective function to efficiently learn our MACB.
  • Actor & Critic Definition. To model the selection of reflections and evaluate the value of the states, we implement a policy network parameterized by MLPs to define our selector’s Actor $\pi_{\bm{{\theta}}}$ and Critic $V_{\bm{\psi}}$, which maps the environmental space $\mathcal{Z}$ to the action space $\mathcal{A}$ and reward function $R$ respectively:
    (7) $\mathrm{Actor\ }\pi_{\bm{\theta}}(\bm{z}):\bm{a}$ $=\mathrm{softmax}(\bm{\theta}\cdot\bm{z})$
    $\mathrm{Critic\ }V_{\bm{\psi}}(\bm{z}):r$ $=\bm{\psi}\cdot\bm{z}$
    where $\bm{\theta}$ and $\bm{\psi}$ are the learnable weights of Actor and Critic respectively.
  • Clip Objective. During the training process, the selector’s Actor uses the $\epsilon$-greedy ($\epsilon=0.1$) strategy to explore the environment, with probability $\epsilon$ to take a random action while with probability ($1-\epsilon$) to exploit the learned policy $\pi_{\bm{\theta}}$. We adopt the clip objective function for PPO training to maximize the reward selection as follows:
    (8) $\mathcal{L}=\mathrm{min}\left(\frac{\pi_{\bm{\theta}}}{\pi_{\mathrm{old}}}A(\bm{z,a}),clip(\frac{\pi_{\bm{\theta}}}{\pi_{\mathrm{old}}},1-\delta,1+\delta)A(\bm{z,a})\right),$
    where $A(\bm{z,a})=\mathbb{E}(R(\bm{z},\bm{a}))-V_{\bm{\psi}}(\bm{z})$ is advantage function used to evaluate $\bm{z}$’s cumulative rewards $\mathbb{E}(R(\bm{z},\bm{a}))$ and current value $V_{\bm{\psi}}(\bm{z})$. $\delta$ is clip threshold.
Inference
During inference, we select the reflection memories with the highest confidence from the Actor, exploit the reflection $Ref_{i}$ with the highest $Imp$ score in the memory, and make a recommendation according to Eq. 2 by replacing the prompt $P$ with the reflection $Ref_{i}$.
Methods Arts Games Instruments
HR@5 HR@10 N@5 N@10 HR@5 HR@10 N@5 N@10 HR@5 HR@10 N@5 N@10
Traditional Deep Recommendations Methods
Caser 0.2555 0.3603 0.1823 0.2161 0.3024 0.4397 0.1997 0.2440 0.2931 0.3968 0.2193 0.2527
FDSA 0.2536 0.3670 0.1875 0.2241 0.3073 0.4466 0.2150 0.2597 0.2842 0.3905 0.2157 0.2497
BERT4Rec 0.2402 0.3613 0.1680 0.2072 0.2796 0.4160 0.1861 0.2299 0.2618 0.3789 0.1954 0.2326
GRU4Rec 0.2431 0.3708 0.1722 0.2129 0.3192 0.4644 0.2083 0.2554 0.3021 0.4021 0.2201 0.2520
SASRec 0.3570 0.4280 0.2788 0.3016 0.2915 0.4348 0.1898 0.2369 0.3450 0.3890 0.2755 0.2896
Training-Free LLM-based Methods
LLM4RS 0.1087 0.1277 0.0826 0.0887 0.1976 0.2233 0.1338 0.1421 0.0900 0.1188 0.0624 0.0717
LLMRank(CoT) 0.3337 0.4061 0.2595 0.2828 0.3508 0.4496 0.2607 0.2925 0.2904 0.3709 0.2202 0.2461
Training-Required LLM-based Methods
LC-Rec 0.3768 0.4309 0.3006 0.3181 0.4140 0.4646 0.3426 0.3593 0.3296 0.3920 0.2486 0.2691
BinLLM 0.2422 0.2896 0.1786 0.2126 0.2701 0.3087 0.2119 0.2523 0.2386 0.2799 0.1958 0.2442
Re2LLM 0.4252 0.4938 0.3535 0.3757 0.4180 0.5198 0.3231 0.3559 0.3521 0.4424 0.2787 0.3074
MoRE 0.4376 0.5043 0.3705 0.3922 0.4773 0.5731 0.3647 0.3957 0.3824 0.4674 0.2981 0.3251
Imp. 2.92% 2.13% 4.81% 4.39% 14.19% 10.25% 6.45% 10.13% 8.61% 5.65% 6.96% 5.76%
Table 1. Recommendation performance comparison. The best and the second-best performances are denoted in bold and underlined fonts, respectively. “N@K” is short for “NDCG@K”. The “Imp.” indicates the percentage improvement of MoRE over the best performances from baselines. denotes MoRE performs significantly better than baselines based on two-tailed paired t-test with Bonferroni correction ($p<0.05$).
4. Experiment
In this section, we conduct extensive experiments to answer the following research questions:
  • RQ1: How does MoRE compare to various existing baselines in terms of recommendation performance?
  • RQ2: How effective are the multi-perspective reflections and reflection selection in MoRE?
  • RQ3: For reflection memory maintenance, how effective are the refining and iteration strategies?
  • RQ4: Does MoRE have an advantage in terms of training time overhead and GPU memory usage?
4.1. Experiment Settings
Dataset. We conduct our experiments on three product categories from the Amazon dataset (Ni et al., 2019): “Arts”, “Video”, and “Instruments”. Each item in these categories has attributes that reflect both explicit user preferences (e.g., titles) and implicit preferences (e.g., brands). To provide sequential recommendations, we follow the approach in (Zheng et al., 2024) and filter out unpopular users and items with fewer than five interactions. The statistics of the pre-processed datasets can be found in App. A.1.
Baselines. We adopt three categories of baselines: traditional deep recommendation methods (e.g., SASRec (Kang and McAuley, 2018), etc.), training-free LLM-based methods (e.g., LLM4RS (Dai et al., 2023), etc.), and training-required LLM-based methods (e.g., Re2LLM (Wang et al., 2024), etc.). See App. A.2 for details.
Evaluation Metrics. To evaluate the performance of our MoRE and other baselines, we adopt two widely used metrics, top-$k$ Hit Ratio (HR@$k$) and Normalized Discounted Cumulative Gain (NDCG@$k$), with $k\in\{5,10\}$. Following (Zheng et al., 2024; Zhou et al., 2022), we employ the leave-one-out strategy for the obtaining of training, validation, and test data. Specifically, for each user behavior sequence, the last item is used as the test data, the penultimate item is used as the validation data, and the remaining interaction records are used for training. Following previous work (Wang et al., 2024), we set every user’s candidate set size $|\mathcal{C}_{u}|$ to 50, where we sample 49 negative items (i.e. items which users do not interact with) from the whole item set for each positive item (i.e. the target item). We perform a ranking evaluation over the candidate set. MoRE and all baselines are required to give a top-10 ranked list for each user.
Implementation Details. We use LLaMa-3111https://huggingface.co/gradientai/Meta-Llama-3-8B-bf16.  (Dubey et al., 2024) as the LLM backbone for MoRE and LLM-based baselines. Implementation details can be found in App. A.3.
4.2. Recommendation Performance
To answer RQ1, we compare the SeqRec performance of MoRE against baselines for 10 trials, with the average results presented in Tab. 1. We can observe the following:
(1) MoRE outperforms all other baselines across each metric on each dataset. On all metrics across three datasets, MoRE outperforms the best baseline by average improvements of 3.56%, 10.25%, and 6.75%, respectively. We attribute this to our proposed LLM-based reflection framework which leverages the reflection capabilities of LLMs to understand both users’ explicit and implicit preferences. Additionally, it enables LLMs to grasp the CF knowledge in a straightforward and comprehensible manner, leading to improved performance in SeqRec.
(2) In general, LLM-based methods demonstrate performance comparable to the traditional methods, which validates the potential of LLMs in SeqRec. Additionally, LLMRank with CoT (Feng et al., 2024) significantly outperforms LLM4RS, indicating that prompt engineering strategies like CoT (Wei et al., 2022) can substantially enhance performance without fine-tuning.
(3) Typically, trained methods demonstrate superior performance compared to their untrained counterparts. Surprisingly, BinLLM delivers relatively underwhelming outcomes. Designed to leverage binary sequences in prompts to capture the correlation between a user-item pair, BinLLM struggles with the nature of SeqRec, which involves a user and multiple candidate items. The insertion of such sequences leads to excessively long prompts, hindering LLMs’ capacity to accurately identify user-item associations.
In summary, MoRE exhibits superior performance in SeqRec, thereby validating the efficacy of our approach.
4.3. Analysis of Multi-Perspective Reflection and Selection
To validate the effectiveness of each perspective of reflection and answer RQ2, we conduct an analytical experiment on Amazon Arts. The results are presented in Tab. 3.
We find that MoRE, equipped with the reflection selection, outperforms alternatives, demonstrating the effectiveness of the multi-perspective reflections and the corresponding perspective selection. Specifically, MoRE outperforms the variants with not only the single reflection perspective but also the dual or triple perspectives simultaneously. This underscores that no single perspective of reflection is universally optimal for all users, thereby confirming the necessity for selective choice among perspectives.
Moreover, MoRE’s selection strategy surpasses both the Random and Greedy selection approaches which respectively denote a random reflection selection from the three memory banks and the selection solely based on downstream recommendation performance. This highlights the effectiveness of our contextual bandit modeling, enabling MoRE to judiciously choose the most suitable reflection for each user.
We also conduct a case study in App. A.5 to further assess the reflection selection of MoRE in comparison to other alternative selection strategies.
(a) Arts
(b) Games
(c) Instruments
Figure 3. Impact of refining and iteration strategies on recommendation performance. We repeated the experiment 5 times, with the mean value for each method represented by the line and the shaded area indicating the standard deviation.
Methods Arts Games Instruments
Time/h GRAM/GB Time/h GRAM/GB Time/h GRAM/GB
LC-Rec (Fine-Tune) 20.382(±0.670) 206.37 21.361(±0.678) 201.35 13.720(±0.436) 199.93
BinLLM (Fine-Tune) 10.209(±0.090) 124.67 11.468(±0.072) 120.57 10.435(±0.094) 115.48
Re2LLM (RL) 4.007(±0.007) 2.77 3.894(±0.005) 1.99 4.152(±0.008) 1.13
MoRE (RL) 3.982(±0.001) 2.77 3.854(±0.002) 1.99 4.093(±0.003) 1.13
Table 2. Training cost comparisons among LLM-based methods on four A800 GPUs. ”Time/h” represents the training time, shown as the mean and standard deviation (SD). ”GRAM/GB” indicates the maximum memory consumption.
Methods N@5 N@10
MoRE 0.3705 0.3922
- Random 0.3421 0.3701
- Greedy 0.3528 0.3786
- $Ref_{\mathrm{EP}}$ only 0.3670 0.3881
- $Ref_{\mathrm{IP}}$ only 0.3571 0.3788
- $Ref_{\mathrm{CF}}$ only 0.3572 0.3788
- $Ref_{\mathrm{EP}}+Ref_{\mathrm{IP}}$ 0.3443 0.3651
- $Ref_{\mathrm{EP}}+Ref_{\mathrm{CF}}$ 0.3311 0.3533
- $Ref_{\mathrm{IP}}+Ref_{\mathrm{CF}}$ 0.3033 0.3217
- $Ref_{\mathrm{EP}}+Ref_{\mathrm{IP}}+Ref_{\mathrm{CF}}$ 0.3051 0.3268
Table 3. Ablation studies on Amazon Arts. “Random” indicates a random reflection selection from the three memory banks, while “Greedy” denotes that reflection selection is based solely on downstream SeqRec on the validation set.
4.4. Analysis of the Refining and Iteration
To answer RQ3, we test the impact of three refining strategies on SeqRec under different iteration rounds, taking reflections on explicit user preferences as an example. From the Fig. 3, we observed that:
(1) Iteration: Overall, as the number of iterations increases, the recommendation performance gradually improves. This indicates that iteration can enhance the quality of reflections, thereby improving the recommendation.
(2) Refining: Refining at group and individual levels shows a significantly better trend over that at the global level during iterations, particularly on Arts (Fig. 3(a)). This improvement may be due to the global-level reflections introducing noise from other users, whereas group- and individual-level demonstrations more effectively elicit personalized reflective capabilities from the reflectors.
Additionally, we found that the refining and iteration strategies are most effective for reflections on explicit preferences, compared to reflections on implicit preferences and collaborative filtering. Detailed experiments and discussions are presented in App. A.4.
4.5. Analysis of LLM Training Cost
To address RQ4, we conduct 5 repeat experiments to compare the training time and maximum memory consumption associated with all the LLM-based baselines that require training. To mitigate the effects of experimental randomness, we perform 5 repetitions of the experiments and report the results as the mean and standard deviation (SD), as shown in Tab. 2. Our findings are as follows:
(1) The training cost of reflection-based methods is lower than fine-tuning methods. This is because the reflection-based methods do not require fine-tuning, thus avoiding significant training time and memory costs.
(2) Under reflection-based setup, MoRE has lower training time and memory consumption compared to Re2LLM. We attribute this to the larger action space of Re2LLM. Specifically, MoRE’s Meta-Reflector does not require selecting specific reflections but only needs to choose among three directions of reflection. In contrast, Re2LLM must first acquire a knowledge base of reflection and, once fixed, conduct PPO training of the retrieval agent within that knowledge base. This means it selects specific reflections from the entire knowledge base, with a larger action space.
In summary, MoRE has the lowest training costs including time overhead and GPU memory usage among all the LLM-based baselines, demonstrating its training efficiency.
5. Conclusion
In this paper, we propose a mixture of reflectors framework, namely MoRE, for modeling and learning the dynamic user preferences in SeqRec. We first introduce three reflectors for each user to generate LLMs’ reflections from the perspectives of explicit user preferences, implicit user preferences, and collaborative signals. Building on these reflectors, we introduce a meta-reflector that evaluates and updates the generated reflections using a self-improving strategy. It selects the most appropriate perspective and corresponding reflections for each user’s current recommendation using a contextual bandit algorithm. Extensive experiments conducted on three benchmark datasets demonstrate that MoRE consistently outperforms traditional recommendation methods and LLM-based recommendation methods with minimal GPU memory usage and training time overhead.
Appendix A Appendix: Experiment Details
A.1. Datasets Statistics
Arts Games Instruments
# Users 55970 55145 27404
# Items 22612 17287 10450
Avg. len 8.80 9.01 8.41
Sparsity 99.96% 99.95% 99.92%
Table 4. Statistics of the 3 pre-processed datasets. Avg. len represents the average length of interacted items per user.
The processed data statistics are summarized in Table 4. Note that “Arts”, “Games” and “Instruments” represent the abbreviations of “Arts, Crafts and Sewing”, “Video Games” and “Musical Instruments” respectively in the Amazon dataset (Amazon Review Data (2018)), where all data include reviews in the range May 1996 - Oct 2018 (Ni et al., 2019).
A.2. Baselines Details
We adopt three types of representative sequential recommendation methods as baselines to validate the effectiveness of our MoRE:
Traditional deep learning recommendation models include:
  • Caser (Tang and Wang, 2018) is a CNN-based method that models user-item interactions as sequences to capture both local and global patterns.
  • GRU4Rec (Hidasi et al., 2015) is an RNN-based SeqRec model with GRU modeling the user behavior sequences.
  • BERT4Rec (Sun et al., 2019) adopts a bidirectional Transformer model and incorporates a masked prediction approach to capture the patterns in item sequences.
  • SASRec (Kang and McAuley, 2018) exploits a unidirectional Transformer-based neural network to model the item sequences and predict the next item.
  • FDSA (Zhang et al., 2019) models both item-level and feature-level transition patterns through a dual self-attention mechanism.
Training-free LLM-based methods, i.e., the prompt-based recommendation methods include:
  • LLM4RS (Dai et al., 2023) investigates the potential of LLMs in enhancing recommendation capabilities by aligning it with traditional information retrieval (IR) ranking techniques, specifically point-wise, pair-wise, and list-wise ranking.
  • LLMRank (Hou et al., 2024) leverages LLMs for zero-shot ranking in recommendation tasks, addressing challenges related to interaction order and bias through specialized prompting techniques.
Training-required LLM-based methods, i.e., fine-tuning and reflection-based recommendation methods include:
  • LC-Rec [^2] To ensure a fair comparison, we ensured that each model’s recommendation process for each user was based on the same item candidate set of size 50. However, since LC-Rec’s training is based on the full item set as the shared candidate set, it is challenging to guarantee that the model does not output items outside the size 50 candidate set for a user solely through the prompt. To address this, we implemented output constraints for the generation process of LC-Rec, ensuring that it outputs only the permitted 50 items for each user (Zheng et al., 2024) is a fine-tuning LLMREC method that incorporates collaborative filtering information by expanding the vocabulary space with special token embeddings. It is currently the state-of-the-art method for LLMREC in the full-candidate SeqRec setting.
  • BinLLM [^3] BinLLM introduces collaborative filtering knowledge to LLMs from the perspective of input text. However, it can only predict the relationship between one user and one item at a time, necessitating multiple inferences to provide recommendations for a user. To ensure a fair comparison, we modified the format of its prompt and reproduced it to make it output a list. (Zhang et al., 2024) is a fine-tuning LLMREC method that incorporates collaborative filtering information by injecting user and item embeddings as binary text sequences into the LLM’s input, without modifying the model’s vocabulary space. It is currently the state-of-the-art method for LLMREC conditioned on the given user-item pair.
  • Re2LLM (Wang et al., 2024) is a reflection-based LLMREC method. It first constructs a reflection knowledge base (KB) with domain knowledge, and then retrieves suitable text in the fixed KB to insert into the LLM’s input, adapting the LLM to session-based recommendation.
(a) Explicit Preference (EP) Reflect
(b) Implicit Preference (IP) Reflect
(c) Collaborative Filtering (CF) Reflect
Figure 4. Impact of iteration on reflections for explicit preferences, implicit preferences, and collaborative filtering information on recommendation performance. The experiments are conducted on the Amazon Games dataset, utilizing an individual-level refining strategy.
A.3. Implementation Details
Considering the cost and efficiency of LLMs, we perform random sampling from three datasets respectively. The principle for sampling users is to exclude those with excessively long interaction sequences before performing random sampling. For each dataset, we sample 1,000 users. Note that items are kept the same as the pre-processed datasets.
[…]
For our MoRE method, we train a DMF (Xue et al., 2017) model as the source of CF Reflection Expert’s CF score (Sec. 3.3). The embedding size of this DMF model is 64 for both users and items. This DMF model is also used to implement the user clustering mentioned in the Refining process in Sec. 3.4. DMF model provides embedding vectors for users in the three datasets, and these embeddings are used to cluster the users in each dataset. We define 20 clusters for each dataset, with the cluster ID being an integer in $[0,19]$, inclusive. For the Iteration process in Sec. 3.4, we set the total iteration to 15, and impacts of Refining and Iteration can be found in Sec. 4.4.
A.4. Impact of Iteration on Different Perspective Reflections on the Amazon Games Dataset
In Sec. 4.4, using LLM reflections on explicit user preferences as an example, we conducted experiments that demonstrated the effectiveness of refining strategies in enhancing reflection quality and improving recommendations.
In this section, we explore the extent to which the iteration strategy improves the quality of reflections on explicit preferences, implicit preferences, and collaborative filtering information. As can be seen from Fig. 4, we have the following three observations.
(1) In terms of the explicit and implicit user preferences, applying the iteration strategies improve the corresponding reflections. The results further verify the effectiveness of our reflection iteration strategy.
(2) Applying iteration strategies to reflections on explicit preferences leads to greater improvements in recommendation performance compared to applying the same strategies to reflections on implicit preferences. We attribute these differences to the reasoning capabilities of the LLM. It is because understanding explicit user preferences is relatively easy for LLMs, as item titles directly reflect user preferences—such as when two products have highly relevant titles. On the other hand, implicit preferences are conveyed through product attributes, which poses greater challenges for LLMs to reason.
(3) The improvement in the CF reflection through iteration strategies is not significant and shows noticeable fluctuations. We believe that even though we circumvent the issue of CF mathematical computations by introducing a tool model for the LLM, numerical comparisons still remain a limitation for the LLM.
In short, the EP and IP reflections are more suitable for the refining and iteration than the CF reflections.
A.5. Case Study
In Sec. 4.3, we conduct ablation studies to quantitatively verify the effectiveness of our reflection selection. In this section, we further validate it qualitatively through a case study. We illustrate the recommendation performances using reflections generated by MoRE and other strategies, shown in the green and brown boxes on the next page.
The case in the brown box illustrates how the original LLMREC, without reflections, failed to hit the recommendation target. Meanwhile, among the various reflection selection methods, the Greedy strategy, which selects reflections based on their performance on the validation set, chooses the EP perspective, ultimately hitting the target at the 7-th position in the recommendation sequence. The Random strategy, on the other hand, randomly selects the IP perspective, which did not hit the recommendation target.
As shown in the green box, MoRE, with the help of the selector, ultimately chooses the CF perspective, hitting the target at the 1-st position in the recommendation sequence. In this case, we observe that the user’s interaction history shows a clear preference for console and action games, with a specific but subtle interest in purchasing items related to console game brands. This preference was not correctly emphasized in the reflections generated from the EP and IP perspectives. Specifically, the IP reflection focuses too much on past preferences for the same brand (e.g., Nintendo games), overlooking potential shifts in the user’s future brand interests. In contrast, based on the Nintendo brand and action games, the CF reflection uses CF information to infer that the user might be interested in PlayStation action games. This approach accurately captured the user’s needs, leading to a correct recommendation.
In short, in this case study, MoRE selects the most appropriate reflective perspective, which greatly assisted the recommendation process.
Moving forward, we believe that employing a reflection mechanism can effectively bridge the semantic gap between LLMs and CF by allowing LLMs to leverage simple score comparisons rather than struggling with the understanding of embeddings. Invoking a straightforward model as a tool to obtain CF scores circumvents the challenge of LLMs understanding CF semantics, thereby endowing LLMs with CF knowledge. Moreover, utilizing reflection along with calling upon simple models offers a cost-effective solution to mitigate various issues LLMs face when adapting to and understanding specialized domains where they might otherwise struggle.
Case Study: MoRE’s Reflection Selection Prompt: I’ve purchased the following products in the past in order: {Ben 10 Omniverse 2 - Nintendo Wii U;…}. And after last previous recommendation (maybe a failed trial), you give the following reflects:{$Ref_{i}$}. Now there are $|\mathcal{C}_{u}|$ candidate products that I can consider to purchase next:$\mathcal{C}_{u}$}.Please rank these $|\mathcal{C}_{u}|$ products by measuring the possibilities that I would like to purchase next most, according to the given purchasing records and reflects. Please think step by step. Please show me your top 10 ranking results with order numbers. Split your output with line break. You MUST rank the given candidates. You can not generate candidates that are not in the given candidate list.
Answer by MoRE perspective selection, where selected reflection (CF reflection) $Ref_{i}=Ref_{\mathrm{CF}}$ = ‘PREFERENCE: The user tends to prefer gaming-related items with high CF scores, particularly those from PlayStation 4 and Xbox 360 platforms, as well as items with high CF scores from previous interactions. ANALYSIS: Although the actor’s previous recommendations are diverse, some items such as {HORI Fighting Commander 4 Controller for PlayStation 4/3, …}, was not included in the actor’s recommendations. This suggests that the actor may have overlooked strong CF signals in favor of more diverse recommendations. SUGGESTIONS: Prioritize gaming-related items with high CF scores, especially those from PlayStation 4 and Xbox 360 platforms, as well as items with high CF scores from previous interactions. Consider including items with CF scores above 0.5, such as {}, to improve the accuracy of the recommendations.’:
1. Street Fighter V - PlayStation 4 Standard Edition (CF score: 0.83), 2. Mass Effect 2 - PC (CF score: 0.78), 3. Assassin’s Creed Liberation HD [Download] (CF score: 0.76), 4. Metal Gear Solid HD Collection (CF score: 0.75), 5. Call of Duty: Advanced Warfare Atlas Limited Edition - Xbox One (CF score: 0.74), 6. Spec Ops: The Line - Playstation 3 (CF score: 0.73), 7. Cities: Skylines [Online Game Code] (CF score: 0.72), 8. Final Fantasy Crystal Chronicles: Echoes of Time - Nintendo Wii (CF score: 0.71), 9. Fatal Fury: Mark of The Wolves (CF score: 0.7), 10. .hack, Part 3: Outbreak (CF score: 0.69)
Case Study: Other Reflection Selectors Prompt: (same form as former prompt)
Answer by recommendation w/o reflection, where selected reflection (EP reflection) $Ref_{i}$ = ‘’:
Didn’t hit at any position) 1. Final Fantasy Legend, 2. Assassin’s Creed Liberation HD [Download], 3. Spinido iPhone 7 Plus Desk Stand Holder Compatible with All Smartphone with Cases- Gold, 4. Phineas and Ferb Ride Again - Nintendo DS, 5. Shadow Warrior, 6. Saints Row IV: National Treasure - Xbox 360, 7. Cabela’s African Adventures - Wii, 8. The Sims 2: Double Deluxe - PC, 9. Disney INFINITY (2.0 Edition) [Download], 10. Metal Gear Solid HD Collection
Answer by Greedy perspective selector, where selected reflection (EP reflection) $Ref_{i}=Ref_{\mathrm{EP}}$ = ‘ PREFERENCE: The user has shown a strong interest in fighting games, particularly those with a classic or retro feel, as well as console-based games. ANALYSIS: The actor’s recommendations did not fully capitalize on the user’s preferences for fighting games and console-based games. The recommended games did not show a strong focus on gameplay and mechanics, and did not prioritize the most engaging and popular titles in the fighting game genre. SUGGESTIONS: Prioritize recommending fighting games, particularly those with a classic or retro feel, and consider suggesting console-based games that cater to the user’s interest in gameplay and mechanics.’:
1. Fatal Fury: Mark of The Wolves, 2. Final Fantasy Crystal Chronicles: Echoes of Time - Nintendo Wii, 3. .hack, Part 3: Outbreak, 4. Capcom Classics Collection, 5. Bad News Baseball - Nintendo NES, 6. Teenage Mutant Ninja Turtles IV: Turtles in Time,
7. Street Fighter V - PlayStation 4 Standard Edition, 8. Metal Gear Solid HD Collection, 9. Final Fantasy Legend, 10. Urban Champion
Answer by Random perspective selector, where selected reflection (IP reflection) $Ref_{i}=Ref_{\mathrm{IP}}$ = ‘ PREFERENCE: Based on the user’s interaction history, it appears that they have a strong preference for games from Nintendo and PlayStation platforms, particularly games from the Wii and Wii U console platforms. The user also shows interest in games with multiplayer features, as evidenced by the inclusion of games like {}. ANALYSIS: To improve the actor’s recommendations, I suggest prioritizing games from Nintendo and PlayStation platforms, particularly games from the Wii and Wii U console platforms. The actor should also focus on games with multiplayer features, as this seems to be an area of interest for the user. Additionally, consider recommending games that offer unique gameplay mechanics, such as {}, which the user has interacted with previously.’:
Didn’t hit at any position) 1. Final Fantasy Crystal Chronicles: Echoes of Time - Nintendo Wii, 2. Mike Tyson’s Punch-Out!! - Nintendo NES, 3. Bad News Baseball - Nintendo NES, 4. Urban Champion - Nintendo NES, 5. Super Mario 64 DS - Nintendo DS, 6. Moshi Monsters - Nintendo DS, 7. LEGO Pirates of the Caribbean - Nintendo DS, 8. Phineas and Ferb Ride Again - Nintendo DS, 9. Cabela’s African Adventures - Wii, 10. Disney Magical World - Nintendo 3DS
Appendix B Appendix: Prompt design
B.1. Base Recommendation Prompt
We provide the base recommendation prompt as follows.
Base recommendation prompt $P$ I’ve purchased the following products in the past in order:$\mathcal{S}_{u}$}. Now there are $|\mathcal{C}_{u}|$ candidate products that I can consider to purchase next: $\mathcal{C}_{u}$}.
Please rank these $|\mathcal{C}_{u}|$ products by measuring the possibilities that I would like to purchase next most, according to the given purchasing records. Please think step by step. Please show me your top 10 ranking results with order numbers. Split your output with line break. You MUST rank the given candidates. You can not generate candidates that are not in the given candidate list.
where $\mathcal{S}_{u}$ is user $u$’s previous interaction sequence in chronological order, $\mathcal{C}_{u}$ is the candidate item set for user $u$, and $|\mathcal{C}_{u}|$ is the candidate item set’s size.
B.2. Recommendation with reflection’s prompt
We provide the prompt of reflection-based recommendation as follows.
Recommendation prompt with reflection $P_{\mathrm{REF}}$ I’ve purchased the following products in the past in order:$\mathcal{S}_{u}$}.
And after last previous recommendation (maybe a failed trial), you give the following reflects:
{$Ref_{i}$}.
Now there are $|\mathcal{C}_{u}|$ candidate products that I can consider to purchase next:
$\mathcal{C}_{u}$}. Please rank these $|\mathcal{C}_{u}|$ products by measuring the possibilities that I would like to purchase next most, according to the given purchasing records. Please think step by step. Please show me your top 10 ranking results with order numbers. Split your output with line break. You MUST rank the given candidates. You can not generate candidates that are not in the given candidate list.
where $\mathcal{S}_{u}$ is user $u$’s previous interaction sequence in chronological order, $\mathcal{C}_{u}$ is the candidate item set for user $u$, and $|\mathcal{C}_{u}|$ is the candidate item set’s size. $Ref_{i}$ is the text of reflection from a specific perspective, including explicit preference (EP) reflection $Ref_{\mathrm{EP}}$, implicit preference (IP) reflection $Ref_{\mathrm{IP}}$ and collaborative filtering (CF) reflection $Ref_{\mathrm{CF}}$.
B.3. Explicit Preference (EP) Reflector’s Generation Prompt
In this section, we provide the detailed information about explicit preference (EP) reflector’s generation prompt, including zero-shot (EP reflection prompt $P_{\mathrm{EP}}$) and few-shot (EP reflection prompt $P_{\mathrm{EP}}$: With demonstrations) version. Here, we first show the zero-shot version, which is used for the beginning of the Refining and Iteration.
EP reflection prompt $P_{\mathrm{EP}}$ You are an expert of giving reflections to the sequential recommender actor. The actor is a sequential recommender. It receives instructions and generates item recommendations based on user’s historical interactions. As an expert, you should provide reflections on the actor’s previous recommendations to help improve actor’s next recommendations for this user.
You need to take it step by step. Step 1: Output 1 or 2 sentences starting with ’PREFERENCE:’.By analyzing the user’s interaction history based on text information such as item titles and categories with your knowledge. (paying more attention to later items in the history), you can predict the user preferences (e.g., trends, anomalies, interest). Step 2. Output 1 or 2 sentences starting with ’ANALYSIS:’. Then analyze the problems with the 10 items recommended by the actor based on user history and preferences. Step 3. Output 1 or 2 sentences starting with ’SUGGESTIONS:’. Finally provide suggestions for the actor to make next recommendations for this user (e.g. prioritize [specific aspect or feature]).
Here are the actor’s recommendations you need to reflect on now (Don’t output any specific titles of items):
Previous instruction:
{$P$}
Standard answer:$\hat{v}$} Actor’s recommendations:$\hat{\mathcal{O}}_{u}$}
Please provide your response based on Previous recommendations below:
where $P$ stands for the previous instruction, also known as the base recommendation prompt (see in App. B.1), which includes user’s previous interaction sequence and candidate item set. All items are displayed with their item titles, which represents the explicit preference of user. $\hat{v}$ is the standard answer of $P$, also known as the ground truth item which user will interacted with in the next time step. $\hat{\mathcal{O}}_{u}$ is actor’s recommendations, also known as the previous predicted item list including each item’s title. The task of $P_{\mathrm{EP}}$ is to generate the EP reflection $Ref_{\mathrm{EP}}$ based on all information mentioned above, especially the item titles which conveys the explicit preference. Then, we display the few-shot version, which is used for bettering Refining and Iteration’s performance by giving valuable demonstrations through specific strategies.
EP reflection prompt $P_{\mathrm{EP}}$: With demonstrations You are an expert of giving reflections to the sequential recommender actor … Finally provide suggestions for the actor to make next recommendations for this user (e.g. prioritize [specific aspect or feature]).
Here are some Examples:
[Example 1 start]
Previous instruction:
{I’ve purchased the following products in the past in order: {Ancient Secrets: Quest for the Golden Key, … , Big City Adventure: Tokyo [Download]}. Now there are 50 candidate products that I can consider to purchase next: {SADES 7.1 Surround Sound Pro … White (SA903), The Jungle Book, … , Left 4 Dead - Game of the Year Edition -Xbox 360}. … Please show me your top 10 ranking results with order numbers. … }
Standard answer:{Jewel Quest Mysteries: The Oracle of Ur [Download]} Actor’s recommendations:{[1. Dark Dimensions: Wax Beauty - PC, 2. Jewel Quest Mysteries: The Oracle of Ur [Download], … , 10. Dungeon Keeper 2 (Jewel Case) - PC]} Please provide your response based on Previous recommendations below:{”PREFERENCE: The user’s interest lies in games that involve puzzle-solving, adventure, and mystery, possibly with a focus on hidden object games, as they have interacted with similar games in the past. ANALYSIS: The actor’s recommendations did not fully capture the user’s preferences, instead suggesting a diverse mix of games across various platforms and genres. While the top 2 recommended items, Jewel Quest Mysteries: The Oracle of Ur [Download] and Dark Dimensions: Wax Beauty - PC, align with the user’s preferences, the subsequent recommendations stray away from the user’s interests. SUGGESTIONS: The actor should prioritize recommending games that fit the user’s preferences for puzzle-solving, adventure, and mystery, such as hidden object games or casual puzzle games, and focus on titles that have a similar tone and atmosphere to the user’s previously interacted items.} [Example 1 end]
[Example 2 start]
[Example 2 end]
Here are the actor’s recommendations you need to reflect on … Please provide your response based on Previous recommendations below:
B.4. Implicit Preference (IP) Reflector’s Generation Prompt
In this section, we provide the detailed information about implicit preference (IP) reflector’s generation prompt, including zero-shot (IP reflection prompt $P_{\mathrm{IP}}$) and few-shot (IP reflection prompt $P_{\mathrm{IP}}$: With demonstrations) version. Here, we first show the zero-shot version, which is used for the beginning of the Refining and Iteration.
IP reflection prompt $P_{\mathrm{IP}}$ You are an expert of giving reflections based on recommended items’ attributes to the sequential recommender actor. The actor is a sequential recommender. It receives instructions and generates item recommendations based on user’s historical interactions. As an expert, you should provide reflections based on the actor’s previous recommendations, especially items’ attributes to help improve actor’s next recommendations for this user.
You need to take it step by step. Step 1: Output 1 or 2 sentences starting with ’PREFERENCE:’. By analyzing the user’s interaction history based on both text information such as items’ titles and items’ attributes with your knowledge. (paying more attention to LATER items’ attributes in the history), you can predict the user preferences (e.g., trends, anomalies, interest). Step 2. Output 1 or 2 sentences starting with ’ANALYSIS:’. Then analyze the problems with the 10 items recommended by the actor based on user history and preferences. Step 3. Output 1 or 2 sentences starting with ’SUGGESTIONS:’. Finally provide suggestions for the actor to make next recommendations for this user (e.g. prioritize [specific aspect, feature or attribute]).
Here are the actor’s recommendations you need to reflect on now (Don’t output any specific titles of items):
Previous instruction and items’ attributes (item titles, categories and brands):
{$P$}
Standard answer:$\hat{v}$} Actor’s recommendations:$\hat{\mathcal{O}}_{u}$}
Please provide your response based on Previous recommendations below:
where $P$ stands for the previous instruction, which includes user’s previous interaction sequence and candidate item set. Different from $P_{\mathrm{EP}}$, all items are displayed with their item titles and side information (including categories and brands), which represents the implicit preference of user. $\hat{v}$ is the standard answer of $P$, also known as the ground truth item which user will interacted with in the next time step. $\hat{\mathcal{O}}_{u}$ is actor’s recommendations, also known as the previous predicted item list including each item’s title. The task of $P_{\mathrm{IP}}$ is to generate the IP reflection $Ref_{\mathrm{IP}}$ based on all information mentioned above, especially the item’s side information which conveys the implicit preference.
Then, we display the few-shot version, which is used for bettering Refining and Iteration’s performance by giving valuable demonstrations through specific strategies.
IP reflection prompt $P_{\mathrm{IP}}$: With demonstrations You are an expert of giving reflections based on recommended items’ attributes to the sequential recommender actor. … Finally provide suggestions for the actor to make next recommendations for this user (e.g. prioritize [specific aspect or feature or attribute]).
Here are some Examples:
[Example 1 start]
Previous instruction and items’ attributes (item titles, categories and brands):{I’ve purchased the following products in the past in order: {{title : “Red Heart Super Saver Yarn, Hunter Green”, categories : “Yarn”, brand : “Red Heart”}, … , {title : “Comfort Grip Crochet Hook 10/Pkg - Assorted Colors (one color per package)”, categories : “Crochet Hooks”, brand : “Kahoot”}}. Now there are 50 candidate products that I can consider to purchase next: {{title : “Liquitex Professional Gloss Gel, Medium, 8 Ounce (5708)”, categories : “Painting, Paints”, brand : “Liquitex”}} … Please show me your top 10 ranking results with order numbers. … }
Standard answer:{Artgraf Watercolor Graphite In Tin}
Actor’s recommendations:{[1. Sizzix 660097 Thinlits Die Magical Butterfly by Pete Hughes, … , 10. addi Knitting Needle Linos Straight 8 inch (20cm) Size US 06 (4.0mm)]}
Please provide your response based on Previous recommendations below: {PREFERENCE: Based on the user’s interaction history, it appears that the user is interested in crafting supplies, particularly related to painting, drawing, and beading. ANALYSIS: The actor’s previous recommendations seem to be a mix of crafting supplies, but they lack a clear focus on the user’s preferences for painting, drawing, and beading. The recommendations also do not prioritize the user’s interaction history, which suggests a strong interest in crafting supplies. SUGGESTIONS: To improve the next recommendations, I suggest prioritizing items that are related to painting, drawing, and beading, as well as considering items that are similar to the user’s previous interactions. The actor should also consider recommending products from brands that are well-known for their quality and relevance to the user’s interests.}
[Example 1 end]
[Example 2 start]
[Example 2 end]
Here are the actor’s recommendations you need to reflect on … Please provide your response based on Previous recommendations below:
Note that implicit preference (side information) are provided as the tuple “{title: …, categories : …, brand: …}” in user’s previous interacted sequence and candidate item set.
B.5. Collaborative Filtering (CF) Reflector’s Generation Prompt
In this section, we provide the detailed information about collaborative filtering (CF) reflector’s generation prompt, including zero-shot (CF reflection prompt $P_{\mathrm{CF}}$) and few-shot (CF reflection prompt $P_{\mathrm{CF}}$: With demonstrations) version. Here, we first show the zero-shot version, which is used for the beginning of the Refining and Iteration.
CF reflection prompt $P_{\mathrm{CF}}$ You are an expert of giving reflections based on user & item’s collaborative filtering scores to the sequential recommender actor. The actor is a sequential recommender. It receives instructions and generates item recommendations based on user’s historical interactions. Then, a collaborative filtering model will provide scores based on current user and generated items. As an expert, you should provide reflections on the actor’s previous recommendations, especially user-item collaborative filtering scores to help improve actor’s next recommendations for this user.
You need to take it step by step. Step 1: Output 1 or 2 sentences starting with ’PREFERENCE:’. By analyzing the user’s interaction history based on both text information such as item titles and categories, and historical items’ collaborative filtering scores with your knowledge. (paying more attention to items with HIGHER collaborative filtering scores in the history),you can predict the user preferences (e.g., trends, anomalies, interest). Step 2. Output 1 or 2 sentences starting with ’ANALYSIS:’. Then analyze the problems with the 0 items recommended by the actor based on user history and preferences. Step 3. Output 1 or 2 sentences starting with ’SUGGESTIONS:’. Finally provide suggestions for the actor to make next recommendations for this user (e.g. prioritize [specific aspect or feature]).
Here are the actor’s recommendations you need to reflect on now (Don’t output any specific titles of items):
Previous instruction and items’ collaborative filtering scores (item titles & CF scores):
{$P$}
Standard answer:$\hat{v}$} Actor’s recommendations:$\hat{\mathcal{O}}_{u}$}
Please provide your response based on Previous recommendations below:
where $P$ stands for the previous instruction, which includes user’s previous interaction sequence and candidate item set. Different from $P_{\mathrm{EP}}$ and $P_{\mathrm{IP}}$, all items are displayed with their item titles and collaborative filtering scores (the CF score, provided by a CF model), which represents the collaborative filtering information of user and items. $\hat{v}$ is the standard answer of $P$, also known as the ground truth item which user will interacted with in the next time step. $\hat{\mathcal{O}}_{u}$ is actor’s recommendations, also known as the previous predicted item list including each item’s title. The task of $P_{\mathrm{CF}}$ is to generate the CF reflection $Ref_{\mathrm{IP}}$ based on all information mentioned above, especially the user-item collaborative filtering information.
CF reflection prompt $P_{\mathrm{CF}}$: With demonstrations You are an expert of giving reflections based on user & item’s collaborative filtering scores to the sequential recommender actor … Finally provide suggestions for the actor to make next recommendations for this user (e.g. prioritize [specific aspect or feature]).
Here are some Examples:
[Example 1 start]
Previous instruction and items’ collaborative filtering scores (item titles & CF scores):{I’ve purchased the following products in the past in order: {{title : “Red Heart Super Saver Yarn, Earth and Sky”, CF score : “0.99847”}, … , {title : “Red Heart Super Saver Yarn, Light Sage”, CF score : “0.99846”}}. Now there are 50 candidate products that I can consider to purchase next: {{title : “Sizzix Bigz Die, 5.5 by 6-Inch, Picture Wheel”, CF score : “0.5”}, … , {title : “RED HEART E302C.0312 Super Saver Jumbo Yarn, Black”, CF score : “0.906811”}}. … Please show me your top 10 ranking results with order numbers. … }
Standard answer:{RED HEART E302C.0312 Super Saver Jumbo Yarn, Black}
Actor’s recommendations:{[1. Red Heart Super Saver Yarn, Earth and Sky, … , 10. Red Heart Super Saver Yarn, Melonberry Print]}
Please provide your response based on Previous recommendations below: {PREFERENCE: Based on the user’s interaction history, it appears that they tend to prefer yarn and fabric related items with high CF scores. The user’s CF scores are highest for Red Heart yarn products, indicating a strong interest in these types of items. ANALYSIS: The actor’s previous recommendations did not fully leverage the user’s preferences, prioritizing items with lower CF scores. The recommended items, although diverse, did not seem to align with the user’s interests in crafts, DIY projects, and art supplies. SUGGESTIONS: To improve the next set of recommendations, I suggest prioritizing items that align with the user’s preferences and CF scores. The actor should consider recommending items related to crafts, DIY projects, and art supplies, such as yarn, fabric, and art materials.}
[Example 1 end]
[Example 2 start]
[Example 2 end]
Here are the actor’s recommendations you need to reflect on … Please provide your response based on Previous recommendations below:
Note that collaborative filtering information are denoted as the tuple “{title: …, CF score: …}” in user’s previous interacted sequence and candidate item set.

### v2

@@
-Enhancing Sequential Recommendations through Multi-Perspective Reflections and Iteration
+MoRE: A Mixture of Reflectors Framework for Large Language Model-Based Sequential Recommendation
 Abstract.
-Sequence recommendation (SeqRec) is a typical task aimed at predicting the next interaction item for users, focusing on understanding user interactive intents and collaborative filtering information.
-Currently, large language models (LLMs) show great potential in recommendation scenarios through prompt-based, fixed reflection library, and fine-tuning approaches.
-However, these approaches face challenges such as lack of supervision, inability to optimize reflection sources, inflexibility to diverse users, and high computational costs.
-Despite achieving encouraging performance, current studies solely exploit reflections on users’ explicit preferences (e.g., item titles).
-They ignore reflecting on implicit preferences and collaborative filtering information, which prevents them from capturing the preference shifts embedded in these signals, neglecting reflections on other aspects of preferences(such as implicit preferences (e.g., brands) and collaborative filtering information.
-As a result, it struggles to capture users’ dynamic preferences.
-Moreover, current methods lack reflection evaluation and iteration, which inevitably affects the quality of reflections and leads to sub-optimal recommendation.
-To address these issues, we propose a Mixture of REflectors framework, namely MoRE, for modeling and learning the dynamic user preferences in sequential recommendations.
-Specifically, we introduce three reflectors for each user to generate LLMs’ reflections on explicit user preferences, implicit user preferences, and collaborative signals.
-Within each reflector, we devise a self-improving strategy called refining-and-iteration to evaluate and update the reflections.
-On top of these reflectors, we introduce a meta-reflector responsible for selecting the most appropriate expert and the corresponding reflections for each user’s current recommendation by applying a contextual bandit algorithm.
-This ensures that we can capture dynamic user preferences effectively.
-Extensive experiments on three real-world datasets demonstrate that our method consistently outperforms state-of-the-art approaches.
-Additionally, we verify that MoRE requires the least training time and GPU memory compared to all other training LLM methods in SeqRec.
-copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06
+Large language models (LLMs) have emerged as a cutting-edge approach in sequential recommendation, leveraging historical interactions to model dynamic user preferences.
+Current methods mainly focus on learning processed recommendation data in the form of sequence-to-sequence text.
+While effective, they exhibit three key limitations: 1) failing to decouple intra-user explicit features (e.g., product titles) from implicit behavioral patterns (e.g., brand loyalty) within interaction histories; 2) underutilizing cross-user collaborative filtering (CF) signals; and 3) relying on inefficient reflection update strategies.
+To address this, We propose MoRE (Mixture of REflectors), which introduces three perspective-aware offline reflection processes to address these gaps.
+This decomposition directly resolves Challenges 1 (explicit/implicit ambiguity) and 2 (CF underutilization).
+Furthermore, MoRE’s meta-reflector employs a self-improving strategy and a dynamic selection mechanism (Challenge 3) to adapt to evolving user preferences.
+First, two intra-user reflectors decouple explicit and implicit patterns from a user’s interaction sequence, mimicking traditional recommender systems’ ability to distinguish surface-level and latent preferences.
+A third cross-user reflector captures CF signals by analyzing user similarity patterns from multiple users’ interactions.
+To optimize reflection quality, MoRE’s meta-reflector employs a offline self-improving strategy that evaluates reflection impacts through comparisons of presence/absence and iterative refinement of old/new versions, with a online contextual bandit mechanism dynamically selecting the optimal perspective for recommendation for each user.
+Experiments on three benchmarks show MoRE outperforms both traditional recommenders and LLM-based methods with minimal computational overhead, validating its effectiveness in bridging LLMs’ semantic understanding with multidimensional recommendation principles.
+Code: https://github.com/E-qin/MoRE-Rec.
+Large Language Model Reflection, Sequential Recommendation
+journalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Ithaca, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Recommender systemsccs: Information systems Language modelsccs: Information systems Personalization
 1. Introduction
 Sequential recommendation (SeqRec), which predicts the next item of interest based on a user’s historical interaction sequence, is crucial for recommender systems.
-The key to this task is to accurately capture users’ dynamic preferences.
+The key to this task is to capture users’ dynamic preferences accurately.
 Large language models (LLMs) are emerging as promising recommenders (LLMREC) due to their vast world knowledge and excellent reasoning abilities.
@@
 Prompt-based methods (Dai et al., 2023; Hou et al., 2023) pre-construct fixed prompts and exploit in-context learning (Dong et al., 2022) to guide the LLM in reasoning the desired item.
-Fine-tuning methods (Zheng et al., 2024; Zhang et al., 2024) inject domain knowledge by fine-tuning an LLM on substantial annotated recommendation data.
+Fine-tuning methods (Zheng et al., 2024; Zhang et al., 2024a; Liao et al., 2024) inject domain knowledge by fine-tuning an LLM on substantial annotated recommendation data.
 Although these methods achieve encouraging performance, the former suffers from an inability to optimize the prompts based on user feedback, while the latter requires substantial computational resources.
-Recently, reflection-based methods (Yao et al., 2023; Wang et al., 2024) have emerged as an alternative approach.
-As shown in Fig.
-1, these methods allow LLMs to reflect on their own outputs based on user history and in turn improve future recommendations.
-Despite their effectiveness, they suffer from the following challenges in modeling and learning user preferences.
-First, these methods typically rely on explicit preferences, such as titles, to represent items and expect that LLMs can capture user interest patterns based on these features.
-However, user interests are often implicitly embedded in the attributes of items.
-For instance, consider a user who is an avid fan of Apple products.
-If she purchases an Apple smartphone and an Apple t-shirt, the connection between these items is based on the brand rather than the item titles.
-The implicit preference, in this case, is the brand “Apple,” which signifies a deeper user preference that is not captured solely by examining the titles of the items.
-Additionally, collaborative filtering (CF) information, which analyzes user behaviors and preferences to find patterns and similarities between users, is not effectively utilized in current approaches.
-This omission leads to a loss of valuable insights that can be gained from user interactions and preferences.
-Second, users’ interests, hidden in their behaviors, are intrinsically dynamic and evolve over time.
-For example, the recommendation at a given time step may depend on reasoning based on explicit preferences, whereas the recommendation at the next step may need to focus on implicit preferences.
-Therefore, it is necessary to dynamically model and select explicit preferences, implicit preferences, and CF information to ensure that the recommender system can adapt to the changing nature of user preferences and provide more accurate and relevant recommendations.
-Third, an ideal reflection update strategy should balance both effectiveness and efficiency, but current methods only manage to achieve one of these qualities.
-One approach (Wang et al., 2024) pre-establishes a fixed pool of reflection candidates to choose from, which may lack the adaptability required for changing user preferences.
-Another approach (Yao et al., 2023) updates the reflection candidate set using large amounts of data.
-While this method is potentially more effective in capturing dynamic interests, it is resource-intensive and less efficient.
 Figure 1.
 Workflow of reflection-based methods for SeqRec.
-LLM reflections are generated offline based on the user’s interaction history ①.
-These reflections are then collected ② and used as prompts for future online recommendations ③.
-Facing the above challenges, we propose a Mixture of REflectors framework, namely MoRE, for modeling and learning the dynamic user preferences in SeqRec.
-Specifically, we introduce three reflectors for each user to collect LLMs’ reflections based on user interaction history, each focusing on different aspects: explicit user preferences, implicit user preferences, and collaborative signals.
-To dynamically update the generated reflections, we devise a meta-reflector with a self-improving strategy called refining and iteration.
-In the refining step, we evaluate the quality of the reflections by measuring the relative improvement in LLMREC, comparing scenarios with and without reflection prompts.
-When the improvement exceeds a certain threshold, the reflection is considered effective.
-In the iteration step, we feed the selected effective reflections into LLMs to generate improved reflections, which can then be sent back to the refining step for a new round of evaluation.
-As such, this offline, self-improving property ensures the effectiveness and efficiency of the reflection update.
-Moreover, the meta-reflector operates on top of three reflectors, responsible for selecting the most appropriate reflection for each user’s current recommendation.
-The selection process is formulated as a contextual bandit problem (Li et al., 2010), which can be efficiently optimized.
-Experimental results demonstrate that MoRE consistently outperforms traditional recommendation methods and LLMREC methods including prompt-based, fine-tuning-based, and reflection-based with minimal GPU memory usage and training time overhead.
+LLM reflections are generated offline based on interaction and recommendation history.
+These reflections are then collected as hints to improve online recommendations.
+Recently, (Yao et al., 2023; Wang et al., 2024) propose reflection-based methods where LLMs refine recommendations by analyzing prediction against user interaction (Fig.
+1).
+While effective, they face three key challenges:
+(1) Ambiguous separation of explicit features and implicit patterns within one single user’s interaction history.
+Effectively modeling user preferences in SeqRec requires distinguishing between explicit and implicit signals within user interaction histories.
+Explicit features are reflected in directly observable item features, such as titles, while implicit patterns are inferred from latent factors in user behavior, such as brand loyalty.
+Current reflection-based recommendation methods (Yao et al., 2023; Wang et al., 2024) predominantly rely on explicit item features to model user interests, assuming that LLMs can effectively capture preference patterns from surface-level semantics alone.
+However, this reliance lacks analysis or reflection on the deeper, implicit connections that often drive user behavior.
+As illustrated in Fig.
+2, consider a user who has interacted with several electronic products and later purchases a T-shirt.
+If the recommender focuses solely on explicit item titles, it struggles to analyze the subtle connection between these interactions and fails to adapt to evolving user interests.
+In this case, the underlying implicit preference is the brand “Apple”, reflecting a deeper pattern that explicit features alone cannot reveal.
+(2) Underutilized collaborative filtering (CF) signals across multiple user interaction histories.
+CF leverages behavioral similarities across users to detect shared interests (e.g., Apple product fans).
+Current approaches (Yao et al., 2023; Wang et al., 2024) neglect such signals, limiting their ability to exploit collective behavioral patterns for improved recommendations.
+(3) Suboptimal reflection update strategies.
+Existing reflection-based methods (Wang et al., 2024; Yao et al., 2023) mainly maintain a fixed reflection pool, lacking personalized dynamic maintenance based on user interactions.
+This inflexibility prevents them from adapting to users’ dynamic preferences.
+Facing these challenges, we propose MoRE (Mixture of REflectors) to address the limitations through three coordinated innovations.
+First, on the offline side, we design specialized reflectors to model both intra-user patterns and cross-user patterns: 1) Two intra-user perspectives disentangle explicit features (e.g., product titles) and implicit behavioral patterns (e.g., brand loyalty) from a single user’s interaction history; 2) A cross-user perspective captures collaborative filtering (CF) signals through item-centric analysis of user-item-user similarity relationships.
+This decomposition directly resolves Challenges 1 (explicit/implicit ambiguity) and 2 (CF underutilization).
+Second, also on the offline side, our self-improving meta-reflector implements a lightweight two-phase update strategy: evaluating reflection quality via LLMREC performance gains (both presence/absence and old/new reflections), then iteratively generating improved reflections using top-performing candidates in an offline loop.
+This approach overcomes Challenge 3 (suboptimal updates) while maintaining computational efficiency.
+Third, on the online side, the meta-reflector dynamically selects optimal reflections per user through contextual bandit optimization (Li et al., 2010).
+Experiments demonstrate MoRE’s superiority over traditional methods and LLMREC variants (prompt-based, fine-tuning-based, and reflection-based baselines) while maintaining low GPU memory and time costs.
 In summary, our contributions are as follows:
-  • We propose an effective and efficient framework, MoRE, for modeling and learning dynamic user preferences in LLM-based SeqRec.
-  • MoRE incorporates three reflectors, each collecting reflections from explicit user preferences, implicit user preferences, and CF signals.
+  • We first propose a dynamic LLM reflection framework is proposed for sequential recommendation to model and learn dynamic user preferences.
+  • MoRE incorporates three reflectors, each collecting reflections from explicit features, implicit patterns, and CF signals.
 It also integrates a meta-reflector, which updates the reflections with a refining-and-iteration strategy and selects the appropriate reflection for LLMREC using a contextual bandit algorithm.
@@
 2.1. Sequential Recommendation
-SeqRec aims to predict a user’s next item of interest by modeling their past interactions chronologically.
-The core challenge lies in accurately capturing users’ dynamic preferences.
-FPMC (Rendle et al., 2010) is an early attempt that addressed this by combining matrix factorization with a Markov chain to capture user preferences.
-In recent years, deep learning methods have significantly advanced recommendation performance.
-For instance, Hidasi et al.
-(2015); Tang and Wang (2018); Li et al.
-(2017) employ convolutional and recurrent neural networks to model local and global interaction patterns.
-Meanwhile, Li et al.
+Figure 2. Examples of LLM reflections from explicit, implicit, and CF perspectives based on user interactions.
+SeqRec predicts users’ next interests by modeling historical interactions chronologically, with the key challenge in capturing dynamic preferences (Shen et al., 2023, 2024; Zhang et al., 2024d, b, 2025; Shi et al., 2024; Zhang et al., 2024c; Tang et al., 2025).
+Recent advances (Hidasi et al., 2015; Tang and Wang, 2018; Li et al., 2017) employ convolutional and recurrent neural networks, while Li et al.
 (2022, 2023); Zhou et al.
-(2022) develop sequential recommendation models based solely on multi-layer perceptrons, demonstrating comparable performance.
-Additionally, Transformer-based models proposed by Kang and McAuley (2018); Sun et al.
+(2022) demonstrate MLP-based approaches achieve competitive performance.
+Kang and McAuley (2018); Sun et al.
 (2019); He et al.
-(2021, 2022) offer precise item-to-item relevance modeling through self-attention, achieving state-of-the-art results.
-Despite their success, these methods often struggle to fully capture the complexity and fluidity of user preferences.
-This limitation is due to the lack of world knowledge and comprehensive reasoning ability.
-Consequently, methods based on large language models have emerged as a promising option for sequential recommendation.
-2.2. Large Language Models for Recommendation
+(2021, 2022) utilize transformer-based models for item relevance modeling, achieving SOTA results.
+Despite progress, existing methods remain limited in holistic preference modeling due to insufficient world knowledge and reasoning capabilities.
+This has spurred recent exploration of LLM-based approaches for sequential recommendation.
+Figure 3.
+The overall architecture of MoRE framework.
+MoRE incorporates multi-perspective reflectors to generate reflections offline ① and a meta-reflector to self-improve these reflections ② and learn to select the most appropriate one ③.
+During online recommendation ④, the meta-reflector selects suitable reflections for the current user.
+2.2. LLM-based Recommendation
 Current approaches to adapting LLMs for SeqRec predominantly fall into two categories: prompt-based methods and fine-tuning methods.
 Prompt-based methods involve manually constructing fixed prompts to guide LLMs in reasoning toward the desired items.
-For example, LLM4RS (Dai et al., 2023) enhances LLM’s recommendation capabilities by customizing prompts to align with traditional ranking, while LLMRank (Hou et al., 2024) leverages LLMs for zero-shot ranking in recommendations through specialized prompting strategies.
+For example, LLM4RS (Dai et al., 2023) enhances LLM’s recommendation capabilities by customizing prompts to align with traditional ranking,
+while LLMRank (Hou et al., 2024) leverages LLMs for zero-shot ranking in recommendations through specialized prompting strategies.
 However, these fixed prompts rely heavily on human expertise and struggle to adapt to diverse users, often resulting in sub-optimal performance (Dong et al., 2022; Wang et al., 2024).
-Fine-tuning methods (Geng et al., 2022), on the other hand, formulate SeqRec as a question answering task and fine-tune accordingly.
-To incorporate recommendation knowledge, they either add special tokens (e.g., Llara (Liao et al., 2024) and LC-Rec (Zheng et al., 2024)) or insert embeddings as texts into the prompt (e.g., BinLLM (Zhang et al., 2024)).
+Fine-tuning methods (Geng et al., 2022; Cui et al., 2022; Zhang et al., 2023b; Bao et al., 2023; Yue et al., 2023; Liao et al., 2024; Lin et al., 2023; Shi et al., 2025b), on the other hand, formulate SeqRec as a question answering task and fine-tune accordingly.
+To incorporate recommendation knowledge, they either add special tokens (e.g., Llara (Liao et al., 2024), LC-Rec (Zheng et al., 2024)) or insert embeddings as texts into the prompt (e.g., BinLLM (Zhang et al., 2024a)).
 Despite these efforts, a significant gap remains between LLM’s pre-training and recommendation-oriented fine-tuning, which necessitates a substantial amount of tuning data to achieve proper alignment.
+Additionally, (Kim et al., 2024; Zhang et al., 2023a; Zhu et al., 2024; Zheng et al., 2024; Zhang et al., 2024a; Shi et al., 2025a) attempt to integrate collaborative filtering information into LLMs.
+While effective, these methods lack an interpretation and analysis of the CF information, which prevents the integration of CF information with the powerful understanding and reasoning capabilities of LLMs, leading to potential suboptimal performance.
 Recently, Wang et al.
 (2024) propose Re2LLM, a reflection-based approach that leverages LLMs to reflect on interaction history to enhance future recommendations.
-However, this approach maintains a static prompt pool and solely reflects on explicit user preferences.
+However, this approach maintains a static prompt pool without decoupled modeling of explicit and implicit preferences and incorporate CF signals.
 3. MoRE: The Proposed Framework
 3.1. Problem Formulation
-In SeqRec, let $\mathcal{V}=\{v_{1},v_{2},...,v_{|\mathcal{V}|}\}$ denote the item set, $\mathcal{U}=\{u_{1},u_{2},...,u_{|\mathcal{U}|}\}$ denote the user set, and $\mathcal{S}_{u}=[v_{1}^{u},v_{2}^{u},...,v_{t}^{u}]$ denote the interaction sequence in chronological order for user $u\in\mathcal{U}$, where $v_{t}^{u}\in\mathcal{V}$ is the item that $u$ has interacted with at time step $t$.
-The goal of this task is to predict the item that $u$ will interact with at $t+1$ given $\mathcal{S}_{u}$:
-(1) $\hat{v}=\operatorname*{arg\,max}_{v\in\mathcal{V}}\Pr(v_{t+1}^{u}=v\mid\mathcal{S}_{u}).$
-In this study, we focus on the LLM-based recommendation. Following (Dai et al., 2023; Hou et al., 2024), we adopt LLM to directly make the prediction as follows:
-(2) $\hat{\mathcal{O}}_{u}=\mathrm{LLM}_{\mathrm{REC}}(P(u,\mathcal{S}_{u},\mathcal{C}_{u})),\hat{\mathcal{O}}_{u}\subseteq\mathcal{C}_{u},$
-where $P$ denotes the manually crafted recommendation prompt; $\mathcal{C}_{u}$ and $\mathcal{O}_{u}$ respectively denote the candidate and predicted item list for $u$; all the parameters are concatenated in the text form; and $\mathrm{LLM}_{\mathrm{REC}}$, the LLM for recommendation, is frozen.
-Please note that reflections are essentially a form of prompts generated by the LLM based on interaction history.
-Therefore, replacing $P()$ with $Ref()$ in Eq.
-2 denotes the reflection-based recommendation.
+In SeqRec, let $\mathcal{V}=\{v_{1},v_{2},...,v_{|\mathcal{V}|}\}$ denotes the item set, $\mathcal{U}=\{u_{1},u_{2},...,u_{|\mathcal{U}|}\}$ denotes the user set, and $\mathcal{S}_{u}^{t}=[v_{u}^{1},v_{u}^{2},...,v_{u}^{t}]$ denotes the interaction sequence in chronological order for user $u\in\mathcal{U}$ up to time step $t$, where $v_{u}^{t}\in\mathcal{V}$ is the item that $u$ has interacted with at time step $t$.
+The goal of this task is to predict the item that $u$ will interact with at $t+1$ given $\mathcal{S}_{u}^{t}$:
+(1) $\hat{v}=\operatorname*{arg\,max}_{v\in\mathcal{V}}\Pr(v_{u}^{t+1}=v\mid\mathcal{S}_{u}^{t}).$
+In this study, we focus on the LLM-based recommendation. Following (Dai et al., 2023; Hou et al., 2024), we adopt LLM to directly make the prediction in a ranking fashion as follows:
+(2) $\hat{\mathcal{O}}_{u}^{t+1}=\mathrm{LLM}_{\mathrm{REC}}(P_{\mathrm{REC}}(u,\mathcal{S}_{u}^{t},\mathcal{C}_{u}^{t+1})),\hat{\mathcal{O}}_{u}^{t+1}\subseteq\mathcal{C}_{u}^{t+1},$
+where $\mathcal{C}_{u}^{t+1}$ and $\hat{\mathcal{O}}_{u}^{t+1}$ respectively denote the candidate and predicted item list for $u$ at step $t+1$; all the parameters are concatenated in the text form; $\mathrm{LLM}_{\mathrm{REC}}$, the LLM for recommendation, is frozen; and $P_{\mathrm{REC}}$ denotes the manually crafted recommendation prompt.
+More specifically, as illustrated in Fig. 1, we focus on the two-stage reflection-based LLMREC.
+In the offline reflection generation stage, given a user $u$, the interaction sequence $\mathcal{S}_{u}^{t-1}$ up to time step $t-1$ and the candidate list $\mathcal{C}_{u}^{t}$, we utilize LLM to make ranking prediction $\hat{\mathcal{O}}_{u}^{t}$ at step $t$:
+(3) $\hat{\mathcal{O}}_{u}^{t}=\mathrm{LLM}_{\mathrm{REC}}(P_{\mathrm{REC}}(u,\mathcal{S}_{u}^{t-1},\mathcal{C}_{u}^{t}),\hat{\mathcal{O}}_{u}^{t}\subseteq\mathcal{C}_{u}^{t}.$
+The LLM then reflects on the predicted list $\hat{\mathcal{O}}_{u}^{t}$ and the ground-truth item $v_{u}^{t}$ to infer $u$’s preferences:
+(4) $Ref_{i,u}=\mathrm{LLM}_{\mathrm{REF}}(P_{i}(u,\mathcal{S}_{u}^{t-1},\mathcal{C}_{u}^{t},\hat{\mathcal{O}}_{u}^{t},v_{u}^{t})),$
+where $\mathrm{LLM}_{\mathrm{REF}}$ denotes a frozen LLM responsible for generating reflections; the subscript $i\in\{EP,IP,CF\}$ for prompt $P_{i}$ and reflection $Ref_{i,u}$ corresponds to the “Explicit Preference”, “Implicit Preference”, and “Collaborative Filtering”, which will be detailed in the next section.
+In the online recommendation stage, the derived $Ref_{i,u}$ is incorporated into Eq. 2 to enhance future recommendations:
+(5) $\hat{\mathcal{O}}_{u}^{t+1}=\mathrm{LLM}_{\mathrm{REC}}(P_{\mathrm{REC}}(u,\mathcal{S}_{u}^{t},\mathcal{C}_{u}^{t+1},Ref_{i,u})),\ \hat{\mathcal{O}}_{u}^{t+1}\subseteq\mathcal{C}_{u}^{t+1}.$
+The details of recommendation prompt $P_{{REC}}$ are as follows:
+Recommendation prompt $P_{REC}$ You are a recommender to recommend items for a specific user.
+The user interacted with items in the following order: $\langle S_{u}\rangle$.
+Reflections on the past recommendation attempt for this user (if any): $\langle Ref_{i,u}\rangle$.
+There are now $|C_{u}|$ candidate items: $\langle C_{u}\rangle$.
+Please consider the user’s historical interaction sequence (and reflections), select appropriate items from the candidates, and rank them to recommend to the user.
+Think step by step.
+Your recommendations:
 3.2. Overall Framework
 We propose MoRE, a novel reflection-based framework, to model the dynamic preferences in LLM-based SeqRec. As shown in Fig. 3, MoRE consist of two key modules:
-Figure 2.
-The overall architecture of our MoRE framework.
-MoRE incorporates multi-perspective reflectors to generate reflections offline ① and a meta-reflector to self-improve these reflections ② and learn to select the most appropriate one ③.
-During online recommendation ④, the meta-reflector selects the suitable reflections for the current user.
-Multi-Perspective Reflectors.
-As discussed in Sec.
-1, user preferences are often implicitly embedded in the attributes of items and can be inferred through collaborative filtering information.
-Therefore, MoRE incorporates this module to drive LLMs to generate reflections from the perspectives of a user’s explicit preferences, implicit preferences, and collaborative filtering information.
-These derived reflections based on historical interactions are considered as hints to enhance LLMs’ future recommendations.
-The details of this module will be introduced in Sec.
-3.3.
-Meta-Reflector.
+Multi-Perspective Reflectors (offline reflection).
+Motivated by the need to decouple explicit/implicit sequential features from user interaction histories and integrate user-item-user collaborative patterns, MoRE employs three perspective-aware reflection processes:
+(1) Explicit/Implicit Decoupling: Two distinct reflections analyze a single user’s history to disentangle explicit preferences (e.g., item titles and descriptions) and implicit preferences (e.g., attribute-level sequential patterns).
+(2) CF-Aware Integration: A third reflection captures cross-user collaborative signals (e.g., rating trends and user similarity) through item-centric analysis.
+As discussed in Sec. 1, these derived reflections based on historical interactions are hints to enhance LLM’s future recommendations. The details of this module will be introduced in Sec. 3.3.
+Meta-Reflector (offline iteration and online selection).
 This module has two main objectives.
@@
 3.5, respectively.
-3.3. Multi-Perspective Reflections Generation
-As illustrated in the left of Fig.
-2, MoRE aims at exploiting users’ historical interaction to generate reflections from the perspectives of explicit preference, implicit preference, and collaborative signals.
+3.3. Offline: Multi-Perspective Reflections
+As illustrated on the left of Fig.
+3, MoRE aims at exploiting users’ historical interaction to generate reflections from the perspectives of explicit preference, implicit preference, and collaborative signals.
 We achieve this goal through the following reflectors:
-Explicit Preference Reflector infers user $u$’s explicit preferences (e.g., item titles) by driving a LLM to reflect the interacted items $\mathcal{S}_{u}$, the candidate items $\mathcal{C}_{u}$, and the difference between LLM’s prediction $\hat{\mathcal{O}}_{u}$ and ground-truth $\mathcal{O}_{u}$.
-LLM generate these reflections simultaneously and return the overall reflection $Ref_{\mathrm{EP}}$, which can be formulated as:
-(3) $Ref_{\mathrm{EP}}=\mathrm{LLM}_{\mathrm{REF}}(P_{\mathrm{EP}}(u,\mathcal{S}_{u},\mathcal{C}_{u},\hat{\mathcal{O}}_{u},\mathcal{O}_{u})),$
-where $\mathrm{LLM}_{\mathrm{REF}}$ is a frozen LLM specific for reflection generation, and $P_{\mathrm{EP}}$ denotes the prompt for reflecting from the perspective of explicit user preferences, which represents items using their titles.
-The details are shown in App.
-B.3.
-Implicit Preference Reflector is designed to capture the preference embedded in the other item attributes (e.g., item brands) from the interaction sequences $\mathcal{S}_{u}$, the candidate set $\mathcal{C}_{u}$ and the LLM prediction $\hat{\mathcal{O}}_{u}$ and label list $\mathcal{O}_{u}$.
-This reflector functions similarly to the explicit preference reflector and adopts a frozen LLM to generate reflections:
-(4) $Ref_{\mathrm{IP}}=\mathrm{LLM}_{\mathrm{REF}}(P_{\mathrm{IP}}(u,\mathcal{S}_{u},\mathcal{C}_{u},\hat{\mathcal{O}}_{u},\mathcal{O}_{u})),$
-where $P_{\mathrm{IP}}$ denotes the reflection prompt from the perspective of implicit user preferences, representing items using their brand names, as detailed in App. B.4.
-Collaborative Filtering Reflector targets guiding LLM to reflect from the perspectives of collaborative filtering information.
-To incorporate the CF signals, instead of introducing new token embeddings into the vocabulary (Zheng et al., 2024) or inserting special sequences into prompts (Zhang et al., 2024), we adopt a pre-trained CF model to rate the CF scores $\mathcal{M}_{u}^{\mathcal{S}}$ and $\mathcal{M}_{u}^{\mathcal{C}}$ respectively for the interacted items for $\mathcal{S}_{u}$ and the candidate items $\mathcal{C}_{u}$.
-These scores are then involved in the input of the reflector, which is formulated as:
-(5) $Ref_{\mathrm{CF}}=\mathrm{LLM}_{\mathrm{REF}}(P_{\mathrm{CF}}(u,\mathcal{S}_{u},\mathcal{C}_{u},\mathcal{M}_{u}^{\mathcal{S}},\mathcal{M}_{u}^{\mathcal{C}},\hat{\mathcal{O}}_{u},\mathcal{O}_{u})),$
-where $P_{\mathrm{CF}}$ denotes the reflection prompt from the perspective of CF, as detailed in App. B.5.
-We maintain three memory banks for each user to store the generated reflections from the corresponding three perspectives.
+Explicit Preference (EP) Reflector captures user $u$’s explicit preferences by leveraging the LLM to reflect on the user’s interaction history $\mathcal{S}_{u}$, the candidate items $\mathcal{C}_{u}^{t}$, and the discrepancy between the predicted ranking $\hat{\mathcal{O}}_{u}^{t}$ generated from LLM-based recommender and the ground-truth $v_{u}^{t}$.
+The reflection $Ref_{\mathrm{EP},u}$ from the perspective of explicit preference can be formulated as follows:
+(6) $Ref_{EP,u}=\mathrm{LLM}_{\mathrm{REF}}(P_{EP}(u,\mathcal{S}_{u}^{t-1},\mathcal{C}_{u}^{t},\hat{\mathcal{O}}_{u}^{t},v_{u}^{t})),$
+where $P_{{EP}}$ denotes the prompt specifically designed to guide the LLM in reflecting explicit user preferences, defined as follows:
+EP reflection prompt $P_{EP}$ You are a reflector for an LLM-based recommender system, understanding the explicit preferences embodied in the user’s historical sequence and analyzing the areas for improvement in past recommendation attempts to provide reflections for the future.
+Explicit preferences are derived from an analysis of recent tendencies reflected in the sequence of item titles and descriptions within the user’s history.
+You need to:
+1.
+Analyze the history with associated text to identify explicit preferences.
+2.
+Analyze the logic and rationale behind past recommendation attempts.
+3.
+Examine potential shortcomings in the past and provide suggestions for improvement.
+User’s historical sequence with related description: $\langle REP_{EP}(S_{u}^{t-1})\rangle$.
+Candidates: $\langle REP_{EP}(C_{u}^{t})\rangle$.
+Past recommendation attempts: $\langle REP_{EP}(\hat{O}_{u}^{t})\rangle$.
+User new interaction (if any): $\langle REP_{EP}(o_{u}^{t})\rangle$.
+Historical reflection demonstrations (if any): $\langle DEMO\rangle$.
+Your reflection:
+Please note that in EP reflector, item titles with descriptions are used to represent $\mathcal{S}_{u}^{t-1},\mathcal{C}_{u}^{t},\hat{\mathcal{O}}_{u}^{t},$ and $v_{u}^{t}$ through $REP_{EP}()$.
+Implicit Preference (IP) Reflector is designed to capture the preference embedded in the other item attributes (e.g., item brands).
+Similar to the EP reflector, the IP reflector takes the interaction history $\mathcal{S}_{u}$ of user $u$, the candidate item list $\mathcal{C}_{u}$, the predicted ranking list $\hat{\mathcal{O}}_{u}$, and the ground-truth item $v_{u}^{t}$ as input, and produce the reflection through a frozen LLM:
+(7) $Ref_{IP,u}=\mathrm{LLM}_{\mathrm{REF}}(P_{IP}(u,\mathcal{S}_{u}^{t-1},\mathcal{C}_{u}^{t},\hat{\mathcal{O}}_{u}^{t},v_{u}^{t})),$
+where $P_{{IP}}$ denotes the prompt for guiding the LLM in reflecting implicit user preferences, defined as follows:
+IP reflection prompt $P_{IP}$ You are a reflector for an LLM-based recommender system, understand the implicit preferences embodied in the user’s historical sequence and analyze the areas for improvement in past recommendation attempts to provide reflections for future.
+Implicit preferences are reflected through the associations shown in subsequences of item attributes within the user’s history, such as subsequences of brands, styles, functions, features, etc.
+Requirements:
+1. Analyze the logic and rationale behind past recommendation attempts.
+2. Focus on attribute subsequences within history to analyze potential associations and causality.
+3. Examine possible shortcomings in the past and provide suggestions.
+Historical attribute seq: $\langle REP_{IP}(S_{u}^{t-1})\rangle$.
+Candidates: $\langle REP_{IP}(C_{u}^{t})\rangle$.
+Past recommendation attempts: $\langle REP_{IP}(\hat{O}_{u}^{t})\rangle$.
+User new interaction (if any): $\langle REP_{IP}(o_{u}^{t})\rangle$.
+Historical reflection demonstrations (if any): $\langle DEMO\rangle$.
+Your reflections:
+In IP reflector, the attributes of items are used to represent $\mathcal{S}_{u}^{t-1},\mathcal{C}_{u}^{t},\hat{\mathcal{O}}_{u}^{t},$ and $v_{u}^{t}$ through $REP_{IP}()$.
+Collaborative Filtering (CF) Reflector aims to integrate patterns and similarities across user interaction sequences into LLM-based recommender systems.
+Unlike approaches that introduce new token embeddings into the vocabulary (Zheng et al., 2024; Liao et al., 2024) or insert special sequences into prompts (Zhang et al., 2024a), we leverage a pre-trained collaborative filtering model $\mathcal{M}_{u}$ to incorporate CF ratings into the LLM reflection process.
+Specifically, the CF reflector takes as input the target user $u$’s interaction history $\mathcal{S}_{u}^{t-1}$, the candidate item set $\mathcal{C}_{u}^{t}$, the predicted ranking $\hat{\mathcal{O}}_{u}^{t}$, and the ground-truth item $v_{u}^{t}$, and generates reflections using a frozen LLM:
+(8) $Ref_{CF,u}=\mathrm{LLM}_{\mathrm{REF}}(P_{CF}(u,\mathcal{S}_{u}^{t-1},\mathcal{C}_{u}^{t},\hat{\mathcal{O}}_{u}^{t},v_{u}^{t})),$
+where $P_{{CF}}$ is the prompt designed to guide the LLM in reflecting CF signals. The prompt structure is detailed as follows:
+CF reflection prompt $P_{CF}$ You are a reflector for an LLM-based recommender system, utilizing a Collaborative Filtering (CF) model to obtain items’ CF signals and analyzing potential improvement in past recommendation attempts for future suggestions.
+The sequence of historical items’ CF ratings reflects patterns and similarities across user sequences.
+Your step-by-step task:
+1. Analyze historical CF ratings to identify preferences, trends, or interest shifts.
+2. Evaluate past recommendation effectiveness by comparing suggested items with actual user interactions.
+3. Analyze potential shortcomings in past recommendation attempts and provide suggestions for improvement.
+Items presented with CF ratings: $\langle REP_{\mathcal{M}_{u}}(S_{u}^{t-1},C_{u}^{t},\hat{O}_{u}^{t},o_{u}^{t})\rangle$ Effective historical reflection demonstrations (if any): $\langle DEMO\rangle$.
+Your reflection:
+In $P_{{CF}}$, $\mathcal{S}_{u}^{t-1},\mathcal{C}_{u}^{t},\hat{\mathcal{O}}_{u}^{t},$ and $v_{u}^{t}$ are represented by their corresponding CF ratings generated by $\mathcal{M}_{u}$ which is denoted as ¡$REP_{\mathcal{M}_{u}}$¿.
+For each user, we maintain three distinct memory banks, each dedicated to storing the generated reflections from one of the three perspectives (Explicit Preference, Implicit Preference, and Collaborative Filtering respectively).
+Each memory bank for a user stores a collection of past reflections for that user from the specific perspective.
 These memory banks are guided by the meta-reflector to improve the reflections, as will be elaborated in the following section.
-3.4. Reflection Memory Maintenance
+3.4. Offline: Reflection Memory Maintenance
 Considering that LLMs may not always accurately generate reflection results, we need a mechanism to ensure the validity of reflections after the meta-reflector obtains them from the three reflectors.
@@
 The greater the performance improvement, the more effective the reflection.
-If the reflection improvement effect of a reflection exceeds a preset threshold, we regard it as an effective one.
+If the reflection improvement effect of a reflection exceeds a preset threshold $h$, we regard it as an effective one.
 Formally, we formulate the reflection improvement effect as:
-(6) $Imp$ $=\mathrm{Metric}(\hat{\mathcal{O}}_{u}^{\mathrm{REF}})-\mathrm{Metric}(\hat{\mathcal{O}}_{u}^{\mathrm{REC}})$
-$\hat{\mathcal{O}}_{u}^{\mathrm{REF}}$ $=\mathrm{LLM}_{\mathrm{REC}}(Ref_{i}(u,\mathcal{S}_{u},\mathcal{C}_{u})),$
-$\hat{\mathcal{O}}_{u}^{\mathrm{REC}}$ $=\mathrm{LLM}_{\mathrm{REC}}(P(u,\mathcal{S}_{u},\mathcal{C}_{u})),$
-where $Ref_{i}\in\{Ref_{\mathrm{EP}},Ref_{\mathrm{IP}},Ref_{\mathrm{CF}}\}$, $\mathrm{Metric}$ denotes the recommendation metric, e.g., NDCG@10.
+(9) $Imp$ $=\mathrm{Metric}(\hat{\mathcal{O}}_{u}^{\mathrm{REF}})-\mathrm{Metric}(\hat{\mathcal{O}}_{u}^{\mathrm{REC}}),$
+$\hat{\mathcal{O}}_{u}^{\mathrm{REF}}$ $=\mathrm{LLM}_{\mathrm{REC}}(P_{\mathrm{REC}}(u,\mathcal{S}_{u}^{t-1},\mathcal{C}_{u}^{t},Ref_{i,u})),$
+$\hat{\mathcal{O}}_{u}^{\mathrm{REC}}$ $=\mathrm{LLM}_{\mathrm{REC}}(P_{\mathrm{REC}}(u,\mathcal{S}_{u}^{t-1},\mathcal{C}_{u}^{t})),$
+where $Ref_{i,u}\in\{Ref_{\mathrm{EP},u},Ref_{\mathrm{IP},u},Ref_{\mathrm{CF},u}\}$, $\mathrm{Metric}$ denotes the recommendation metric, e.g., NDCG@10.
 Refining. As we filter out reflections that offer less or negative SeqRec improvement using the above measure, we further refine the selected reflections with three strategies:
 (1) Global Level: We select the reflections that bring the greatest average improvement for all users. In this approach, all users share the same reflections.
-(2) Group Level: We use a collaborative filtering (CF) model to construct embeddings for all users and then cluster them using the K-means++ algorithm (Bahmani et al., 2012).
+(2) Group Level: We use a CF model to construct embeddings for all users and then cluster them using the K-means++ algorithm (Bahmani et al., 2012).
 We then choose the reflections that provide the greatest average improvement for users within the same cluster, ensuring that users in the same cluster share the same reflections.
 (3) Individual Level: We select the reflections that bring the greatest improvement for each user. This strategy leads to personalized reflections.
-All three of the aforementioned strategies are greedy, which may lead to local optima.
+All of the three strategies are greedy, which may lead to local optima.
 To mitigate this issue, we propose using the improvement effect brought about by each reflection as the sampling probability.
-Reflections sampled from the filtered set serve as demonstrations and are exploited in the subsequent iteration step.
+Reflections sampled from the filtered set serve as in-context learning demonstrations.
+These demonstrations are then used to task the three reflectors with generating improved reflections in subsequent iteration steps, enabling the LLM to self-improve its reflection quality.
 Iteration.
@@
 We will further validate the impact of reflection iteration on LLM recommendation performance in the experiments.
-3.5. Reflection Selection
+3.5. Online: Reflection Perspective Selection
 The second objective of the meta-reflector is to model preference shifts in SeqRec.
-Since we have designed three memory banks for each user to collect effective reflections from three different perspectives, we require the meta-reflector to select one of these reflections for recommendation to achieve this goal.
+Given three memory banks for each user to collect effective reflections from three different perspectives, we require the meta-reflector to select one of these reflections for recommendation to achieve this goal.
 We formulate the decision-making process as a Multi-Armed Contextual Bandit (MACB) problem (Li et al., 2010), and adopt the Proximal Policy Optimization (PPO) algorithm (Schulman et al., 2017) to efficiently learn to select the reflections.
 Formally, we define MACB as a tuple $(\mathcal{Z},\mathcal{A},\mathcal{R})$, where $\mathcal{Z},\mathcal{A},\mathcal{R}$ denote the state space, action (arm) space, and reward function, respectively.
-  • State $\bm{z}\in\mathcal{Z}$ is defined as the recommendation context.
-We represent it by concatenating the embedding of user $u$ and the interaction sequence $\mathcal{S}_{u}$, i.e., $\bm{z}=[emb_{u},emb_{v}]$, where the embeddings are derived from a CF model, and $emb_{v}$ denotes the average embedding for $\mathcal{S}_{u}$.
-  • Action $\bm{a}\in\mathcal{A}$ is defined as the selection of a reflection from three memory banks.
+(1) state $\bm{z}\in\mathcal{Z}$ represents recommendation context, constructed by concatenating CF-derived embeddings: $\bm{z}=[emb_{u},emb_{v}]$ where $emb_{u}$ is the user embedding and $emb_{v}$ the average interaction embedding from $\mathcal{S}_{u}^{t-1}$.
+(2) Action $\bm{a}\in\mathcal{A}$ is defined as the selection of a reflection from three memory banks.
 We represent $\bm{a}$ as a 3-dimensional vector, with each dimension corresponding to an arm, each arm representing a reflection memory.
-  • Reward is defined to assess the recommendation improvement brought by the reflections.
+(3) Reward is defined to assess the recommendation improvement brought by the reflections.
 We reuse the measure in Eq.(9) and represent the reward as $R(\bm{z,a})=Imp$.
@@
 3.4, $R(\bm{z,a})$ can be both positive and negative, ensuring training efficiency.
-  • Replay Buffer is designed to facilitate efficiency in policy optimization.
+(4) Replay Buffer is designed to facilitate efficiency in policy optimization.
 We represent it as $D=(\bm{z},\bm{a},R(\bm{z},\bm{a}),\bm{z}^{\prime})$ to store the tuples of observed state, action, reward, and next state.
 With the records in the replay buffer, we can refine successful policies and learn from erroneous trials.
+Table 1. Statistics of the 3 pre-processed datasets.
+Dataset #Users #Items Avg. len Sparsity
+Arts 55970 22612 8.80 99.96%
+Games 55145 17287 9.01 99.95%
+Instruments 27404 10450 8.41 99.92%
+Table 2.
+Recommendation performance comparison.
+All LLM-based methods utilize Llama-3 and ensure an identical candidate set for a fair comparison.
+The best and the second-best performances are denoted in bold and underlined fonts, respectively.
+“N@K” is short for “NDCG@K”.
+The “Imp.” indicates the percentage improvement of MoRE over the best performances from baselines.
+“-” means that the baseline can only output a top-1 prediction and cannot rank or recall multiple items.
+denotes MoRE performs significantly better than baselines based on two-tailed paired t-test with Bonferroni correction ($p<0.05$).
+Methods Arts Games Instruments
+HR@1 HR@5 HR@10 N@5 N@10 HR@1 HR@5 HR@10 N@5 N@10 HR@1 HR@5 HR@10 N@5 N@10
+Caser 0.103 0.2555 0.3603 0.1823 0.2161 0.0899 0.3024 0.4397 0.1997 0.2440 0.1349 0.2931 0.3968 0.2193 0.2527
+FDSA 0.1211 0.2536 0.3670 0.1875 0.2241 0.1196 0.3073 0.4466 0.2150 0.2597 0.1403 0.2842 0.3905 0.2157 0.2497
+BERT4Rec 0.0896 0.2402 0.3613 0.1680 0.2072 0.0899 0.2796 0.4160 0.1861 0.2299 0.1233 0.2618 0.3789 0.1954 0.2326
+GRU4Rec 0.1001 0.2431 0.3708 0.1722 0.2129 0.0929 0.3192 0.4644 0.2083 0.2554 0.1340 0.3021 0.4021 0.2201 0.2520
+SASRec 0.1850 0.3570 0.4280 0.2788 0.3016 0.1900 0.2915 0.4348 0.1898 0.2369 0.1832 0.3327 0.3701 0.2594 0.2809
+LLM4RS 0.0543 0.1087 0.1277 0.0826 0.0887 0.0662 0.1976 0.2233 0.1338 0.1421 0.0330 0.0900 0.1188 0.0624 0.0717
+LLMRank(CoT) 0.1725 0.3337 0.4061 0.2595 0.2828 0.1611 0.3508 0.4496 0.2607 0.2925 0.1457 0.2904 0.3709 0.2202 0.2461
+LC-Rec 0.2145 0.3768 0.4309 0.3006 0.3181 0.2599 0.4140 0.4646 0.3426 0.3593 0.1592 0.3296 0.3920 0.2486 0.2691
+Re2LLM 0.2736 0.4252 0.4938 0.3535 0.3757 0.2223 0.4180 0.5198 0.3231 0.3559 0.2011 0.3521 0.4424 0.2787 0.3074
+MoRE 0.2898 0.4376 0.5043 0.3705 0.3922 0.2569 0.4773 0.5731 0.3647 0.3957 0.2163 0.3824 0.4674 0.2981 0.3251
+Imp. 5.92% 2.92% 2.13% 4.81% 4.39% -1.15% 14.19% 10.25% 6.45% 10.13% 7.56% 8.61% 5.65% 6.96% 5.76%
 Training
-We apply the PPO algorithm (Schulman et al., 2017) with an Actor & Critic networks and a Clip objective function to efficiently learn our MACB.
-  • Actor & Critic Definition.
+We apply the PPO algorithm (Schulman et al., 2017) with an Actor & Critic network and a Clip objective function to efficiently learn MACB.
+(1) Actor & Critic Definition.
 To model the selection of reflections and evaluate the value of the states, we implement a policy network parameterized by MLPs to define our selector’s Actor $\pi_{\bm{{\theta}}}$ and Critic $V_{\bm{\psi}}$, which maps the environmental space $\mathcal{Z}$ to the action space $\mathcal{A}$ and reward function $R$ respectively:
-    (7) $\mathrm{Actor\ }\pi_{\bm{\theta}}(\bm{z}):\bm{a}$ $=\mathrm{softmax}(\bm{\theta}\cdot\bm{z})$
-    $\mathrm{Critic\ }V_{\bm{\psi}}(\bm{z}):r$ $=\bm{\psi}\cdot\bm{z}$
-    where $\bm{\theta}$ and $\bm{\psi}$ are the learnable weights of Actor and Critic respectively.
-  • Clip Objective.
+(10) $\mathrm{Actor\ }\pi_{\bm{\theta}}(\bm{z}):\bm{a}$ $=\mathrm{softmax}(\bm{\theta}\cdot\bm{z}),$
+$\mathrm{Critic\ }V_{\bm{\psi}}(\bm{z}):r$ $=\bm{\psi}\cdot\bm{z},$
+where $\bm{\theta}$ and $\bm{\psi}$ are the parameters of Actor and Critic respectively.
+(2) Clip Objective.
 During the training process, the selector’s Actor uses the $\epsilon$-greedy ($\epsilon=0.1$) strategy to explore the environment, with probability $\epsilon$ to take a random action while with probability ($1-\epsilon$) to exploit the learned policy $\pi_{\bm{\theta}}$.
@@
 Inference
-During inference, we select the reflection memories with the highest confidence from the Actor, exploit the reflection $Ref_{i}$ with the highest $Imp$ score in the memory, and make a recommendation according to Eq.
-2 by replacing the prompt $P$ with the reflection $Ref_{i}$.
+During inference, we select the reflection memories with the highest confidence from the Actor, exploit the reflection $Ref_{i,u}$ with the highest $Imp$ score in the memory, and make a recommendation according to Eq.(2) using the prompt $P_{\mathrm{REC}}$.
+4. Experiment
+In this section, we conduct extensive experiments to answer the following research questions: RQ1: How does MoRE compare to various existing baselines in terms of recommendation performance?
+RQ2: How effective are the multi-perspective reflections and reflection perspective selection in MoRE?
+RQ3: For reflection memory maintenance, how effective are the refining and iteration strategies?
+RQ4: Does MoRE have an advantage in terms of training cost?
+Table 3. Comparison of HR@1 Performance with LLM baselines that inherently only provide top-1 prediction.
 Methods Arts Games Instruments
-HR@5 HR@10 N@5 N@10 HR@5 HR@10 N@5 N@10 HR@5 HR@10 N@5 N@10
-Traditional Deep Recommendations Methods
-Caser 0.2555 0.3603 0.1823 0.2161 0.3024 0.4397 0.1997 0.2440 0.2931 0.3968 0.2193 0.2527
-FDSA 0.2536 0.3670 0.1875 0.2241 0.3073 0.4466 0.2150 0.2597 0.2842 0.3905 0.2157 0.2497
-BERT4Rec 0.2402 0.3613 0.1680 0.2072 0.2796 0.4160 0.1861 0.2299 0.2618 0.3789 0.1954 0.2326
-GRU4Rec 0.2431 0.3708 0.1722 0.2129 0.3192 0.4644 0.2083 0.2554 0.3021 0.4021 0.2201 0.2520
-SASRec 0.3570 0.4280 0.2788 0.3016 0.2915 0.4348 0.1898 0.2369 0.3450 0.3890 0.2755 0.2896
-Training-Free LLM-based Methods
-LLM4RS 0.1087 0.1277 0.0826 0.0887 0.1976 0.2233 0.1338 0.1421 0.0900 0.1188 0.0624 0.0717
-LLMRank(CoT) 0.3337 0.4061 0.2595 0.2828 0.3508 0.4496 0.2607 0.2925 0.2904 0.3709 0.2202 0.2461
-Training-Required LLM-based Methods
-LC-Rec 0.3768 0.4309 0.3006 0.3181 0.4140 0.4646 0.3426 0.3593 0.3296 0.3920 0.2486 0.2691
-BinLLM 0.2422 0.2896 0.1786 0.2126 0.2701 0.3087 0.2119 0.2523 0.2386 0.2799 0.1958 0.2442
-Re2LLM 0.4252 0.4938 0.3535 0.3757 0.4180 0.5198 0.3231 0.3559 0.3521 0.4424 0.2787 0.3074
-MoRE 0.4376 0.5043 0.3705 0.3922 0.4773 0.5731 0.3647 0.3957 0.3824 0.4674 0.2981 0.3251
-Imp. 2.92% 2.13% 4.81% 4.39% 14.19% 10.25% 6.45% 10.13% 8.61% 5.65% 6.96% 5.76%
+TALLRec 0.2559 0.1975 0.1588
+BinLLM 0.2401 0.2347 0.1875
+LLaRA 0.2803 0.2263 0.1763
+A-LLMREC 0.2193 0.2342 0.1897
+MoRE 0.2898 0.2569 0.2163
 Table 4.
-Recommendation performance comparison.
-The best and the second-best performances are denoted in bold and underlined fonts, respectively.
-“N@K” is short for “NDCG@K”.
-The “Imp.” indicates the percentage improvement of MoRE over the best performances from baselines.
-denotes MoRE performs significantly better than baselines based on two-tailed paired t-test with Bonferroni correction ($p<0.05$).
-4. Experiment
-In this section, we conduct extensive experiments to answer the following research questions:
-  • RQ1: How does MoRE compare to various existing baselines in terms of recommendation performance?
-  • RQ2: How effective are the multi-perspective reflections and reflection selection in MoRE?
-  • RQ3: For reflection memory maintenance, how effective are the refining and iteration strategies?
-  • RQ4: Does MoRE have an advantage in terms of training time overhead and GPU memory usage?
+Ablation studies on Amazon Arts.
+“Random” indicates a random reflection selection from the three memory banks, while “Greedy” denotes that reflection selection is based solely on downstream SeqRec on the validation set.
+$Ref_{\mathrm{all}}$ denotes $Ref_{\mathrm{EP}}+Ref_{\mathrm{IP}}+Ref_{\mathrm{CF}}$.
+Note: “+” denotes concatenation, which may cause conflicting reflections leading to suboptimal performance.
+MoRE achieves the best results through user-personalized RL selection.
+Methods HR@1 HR@5 HR@10 N@5 N@10
+MoRE 0.2898 0.4376 0.5043 0.3705 0.3922
+- Random 0.2736 0.4023 0.4881 0.3421 0.3701
+- Greedy 0.2812 0.4156 0.4957 0.3528 0.3786
+- $Ref_{\mathrm{EP}}$ only 0.2850 0.4356 0.5033 0.3670 0.3881
+- $Ref_{\mathrm{IP}}$ only 0.2850 0.4214 0.4881 0.3571 0.3788
+- $Ref_{\mathrm{CF}}$ only 0.2850 0.4290 0.4871 0.3572 0.3788
+- $Ref_{\mathrm{EP}}+Ref_{\mathrm{IP}}$ 0.2583 0.4071 0.4700 0.3443 0.3651
+- $Ref_{\mathrm{EP}}+Ref_{\mathrm{CF}}$ 0.2707 0.3927 0.4623 0.3311 0.3533
+- $Ref_{\mathrm{IP}}+Ref_{\mathrm{CF}}$ 0.2402 0.3584 0.4156 0.3033 0.3217
+- $Ref_{\mathrm{all}}$ 0.2393 0.3594 0.4271 0.3051 0.3268
 4.1. Experiment Settings
 Dataset.
-We conduct our experiments on three product categories from the Amazon dataset (Ni et al., 2019): “Arts”, “Video”, and “Instruments”.
-Each item in these categories has attributes that reflect both explicit user preferences (e.g., titles) and implicit preferences (e.g., brands).
-To provide sequential recommendations, we follow the approach in (Zheng et al., 2024) and filter out unpopular users and items with fewer than five interactions.
-The statistics of the pre-processed datasets can be found in App.
-A.1.
+Following (Zheng et al., 2024), we conduct our experiments on three subsets of the Amazon dataset (Ni et al., 2019): “Arts”, “Video”, and “Instruments” [^1] Note that “Arts”, “Games” and “Instruments” represent the abbreviations of “Arts, Crafts and Sewing”, “Video Games” and “Musical Instruments” respectively in the Amazon Review dataset (2018)..
+Each item in these subsets contains attributes that capture both explicit user preferences (e.g., item titles) and implicit preferences (e.g., attributes like brands and functional traits ).
+We follow (Zheng et al., 2024) to preprocess the data to achieve a fair comparison, with the statistical information shown in Tab.
+1.
 Baselines.
-We adopt three categories of baselines: traditional deep recommendation methods (e.g., SASRec (Kang and McAuley, 2018), etc.), training-free LLM-based methods (e.g., LLM4RS (Dai et al., 2023), etc.), and training-required LLM-based methods (e.g., Re2LLM (Wang et al., 2024), etc.).
-See App.
-A.2 for details.
+We adopt three types of SeqRec methods as baselines: (1) Traditional Deep Learning Models: Caser (Tang and Wang, 2018) captures local and global patterns via CNN; GRU4Rec (Hidasi et al., 2015) models sequences with GRUs; SASRec (Kang and McAuley, 2018) and BERT4Rec (Sun et al., 2019) employ uni-/bidirectional Transformers for next-item prediction; FDSA (Zhang et al., 2019) models item-feature transitions via dual attention.
+(2) Training-Free LLM Methods: LLM4RS (Dai et al., 2023) aligns LLMs with IR ranking strategies; LLMRank (Hou et al., 2024) achieves zero-shot ranking via specialized prompting.
+(3) Training-Required LLM Methods: LC-Rec (Zheng et al., 2024) expands vocabulary for collaborative filtering; TALLRec (Bao et al., 2023) uses instruction tuning with text descriptions; BinLLM (Zhang et al., 2024a) encodes user-item interactions as binary sequences; Re2LLM (Wang et al., 2024) employs fixed reflection pools for suggestions; LLaRA (Liao et al., 2024) combines user knowledge and behavior patterns via LoRA; A-LLMREC (Kim et al., 2024) aligns CF embeddings with LLM vocabulary.
 Evaluation Metrics.
-To evaluate the performance of our MoRE and other baselines, we adopt two widely used metrics, top-$k$ Hit Ratio (HR@$k$) and Normalized Discounted Cumulative Gain (NDCG@$k$), with $k\in\{5,10\}$.
-Following (Zheng et al., 2024; Zhou et al., 2022), we employ the leave-one-out strategy for the obtaining of training, validation, and test data.
+We adopt two widely used metrics, top-$k$ Hit Ratio (HR@$k$) and Normalized Discounted Cumulative Gain (NDCG@$k$), with $k\in\{5,10\}$.
+We follow (Zheng et al., 2024; Zhou et al., 2022) to employ the leave-one-out strategy for the obtaining of training, validation, and test data.
 Specifically, for each user behavior sequence, the last item is used as the test data, the penultimate item is used as the validation data, and the remaining interaction records are used for training.
-Following previous work (Wang et al., 2024), we set every user’s candidate set size $|\mathcal{C}_{u}|$ to 50, where we sample 49 negative items (i.e.
-items which users do not interact with) from the whole item set for each positive item (i.e.
-the target item).
-We perform a ranking evaluation over the candidate set.
-MoRE and all baselines are required to give a top-10 ranked list for each user.
-Implementation Details.
-We use LLaMa-3111https://huggingface.co/gradientai/Meta-Llama-3-8B-bf16.
-(Dubey et al., 2024) as the LLM backbone for MoRE and LLM-based baselines.
-Implementation details can be found in App.
-A.3.
-4.2. Recommendation Performance
-To answer RQ1, we compare the SeqRec performance of MoRE against baselines for 10 trials, with the average results presented in Tab. 1. We can observe the following:
-(1) MoRE outperforms all other baselines across each metric on each dataset.
-On all metrics across three datasets, MoRE outperforms the best baseline by average improvements of 3.56%, 10.25%, and 6.75%, respectively.
-We attribute this to our proposed LLM-based reflection framework which leverages the reflection capabilities of LLMs to understand both users’ explicit and implicit preferences.
-Additionally, it enables LLMs to grasp the CF knowledge in a straightforward and comprehensible manner, leading to improved performance in SeqRec.
-(2) In general, LLM-based methods demonstrate performance comparable to the traditional methods, which validates the potential of LLMs in SeqRec.
-Additionally, LLMRank with CoT (Feng et al., 2024) significantly outperforms LLM4RS, indicating that prompt engineering strategies like CoT (Wei et al., 2022) can substantially enhance performance without fine-tuning.
-(3) Typically, trained methods demonstrate superior performance compared to their untrained counterparts.
-Surprisingly, BinLLM delivers relatively underwhelming outcomes.
-Designed to leverage binary sequences in prompts to capture the correlation between a user-item pair, BinLLM struggles with the nature of SeqRec, which involves a user and multiple candidate items.
-The insertion of such sequences leads to excessively long prompts, hindering LLMs’ capacity to accurately identify user-item associations.
-In summary, MoRE exhibits superior performance in SeqRec, thereby validating the efficacy of our approach.
-4.3. Analysis of Multi-Perspective Reflection and Selection
-To validate the effectiveness of each perspective of reflection and answer RQ2, we conduct an analytical experiment on Amazon Arts. The results are presented in Tab. 3.
-We find that MoRE, equipped with the reflection selection, outperforms alternatives, demonstrating the effectiveness of the multi-perspective reflections and the corresponding perspective selection.
-Specifically, MoRE outperforms the variants with not only the single reflection perspective but also the dual or triple perspectives simultaneously.
-This underscores that no single perspective of reflection is universally optimal for all users, thereby confirming the necessity for selective choice among perspectives.
-Moreover, MoRE’s selection strategy surpasses both the Random and Greedy selection approaches which respectively denote a random reflection selection from the three memory banks and the selection solely based on downstream recommendation performance.
-This highlights the effectiveness of our contextual bandit modeling, enabling MoRE to judiciously choose the most suitable reflection for each user.
-We also conduct a case study in App. A.5 to further assess the reflection selection of MoRE in comparison to other alternative selection strategies.
+Following (Hou et al., 2024; Wang et al., 2024; Kim et al., 2024; Liao et al., 2024), we randomly sample $|\mathcal{C}_{u}|-1$ negative items to construct $|\mathcal{C}_{u}|$ candidates with 1 target item for each user.
+Following  (Wang et al., 2024), $|\mathcal{C}_{u}|$ is set to 50 for all methods.
+All methods except those inherently limited to top-1 prediction (without ranking) must provide top-10 ranked items.
+[…]
+Table 5.
+Training cost comparisons among LLM-based methods on four A800.
+“Training time” is measured in hours, shown as the mean and standard deviation (SD).
+“VRAM” denotes the Peak VRAM Usage, measured in GB.
+MoRE’s RL framework minimizes computation cost through compact action space ($|\mathcal{A}|=3$, 3 perspectives).
+Methods Arts Games Instruments
+Training time VRAM Training time VRAM Training time VRAM
+TALLRec (Fine-Tune) 26.228(±0.818) 46.98 18.018(±0.290) 43.24 18.338(±0.487) 42.95
+BinLLM (Fine-Tune) 10.209(±0.540) 124.67 11.468(±0.328) 120.57 10.435(±0.451) 115.48
+LLaRA (Fine-Tune) 9.013(±0.161) 77.92 6.405(±0.281) 65.83 6.863(±0.187) 74.36
+A-LLMREC (Fine-Tune) 56.975(±1.621) 217.74 55.975(±1.058) 202.09 56.316(±1.992) 231.63
+LC-Rec (Fine-Tune) 20.382(±0.670) 206.37 21.361(±0.678) 201.35 13.720(±0.436) 199.93
+Re2LLM (RL, $|\mathcal{A}|=20$) 4.007(±0.007) 2.77 3.894(±0.005) 1.99 4.152(±0.008) 1.13
+MoRE (RL, $|\mathcal{A}|=3$) 3.982(±0.001) 2.77 3.854(±0.002) 1.99 4.093(±0.003) 1.13
 (a) Arts
@@
 (c) Instruments
+(d) Iteration for 3 Perspectives
 Figure 4.
@@
 We repeated the experiment 5 times, with the mean value for each method represented by the line and the shaded area indicating the standard deviation.
-Methods Arts Games Instruments
-Time/h GRAM/GB Time/h GRAM/GB Time/h GRAM/GB
-LC-Rec (Fine-Tune) 20.382(±0.670) 206.37 21.361(±0.678) 201.35 13.720(±0.436) 199.93
-BinLLM (Fine-Tune) 10.209(±0.090) 124.67 11.468(±0.072) 120.57 10.435(±0.094) 115.48
-Re2LLM (RL) 4.007(±0.007) 2.77 3.894(±0.005) 1.99 4.152(±0.008) 1.13
-MoRE (RL) 3.982(±0.001) 2.77 3.854(±0.002) 1.99 4.093(±0.003) 1.13
-Table 2.
-Training cost comparisons among LLM-based methods on four A800 GPUs.
-”Time/h” represents the training time, shown as the mean and standard deviation (SD).
-”GRAM/GB” indicates the maximum memory consumption.
-Methods N@5 N@10
-MoRE 0.3705 0.3922
-- Random 0.3421 0.3701
-- Greedy 0.3528 0.3786
-- $Ref_{\mathrm{EP}}$ only 0.3670 0.3881
-- $Ref_{\mathrm{IP}}$ only 0.3571 0.3788
-- $Ref_{\mathrm{CF}}$ only 0.3572 0.3788
-- $Ref_{\mathrm{EP}}+Ref_{\mathrm{IP}}$ 0.3443 0.3651
-- $Ref_{\mathrm{EP}}+Ref_{\mathrm{CF}}$ 0.3311 0.3533
-- $Ref_{\mathrm{IP}}+Ref_{\mathrm{CF}}$ 0.3033 0.3217
-- $Ref_{\mathrm{EP}}+Ref_{\mathrm{IP}}+Ref_{\mathrm{CF}}$ 0.3051 0.3268
-Table 3.
-Ablation studies on Amazon Arts.
-“Random” indicates a random reflection selection from the three memory banks, while “Greedy” denotes that reflection selection is based solely on downstream SeqRec on the validation set.
-4.4. Analysis of the Refining and Iteration
-To answer RQ3, we test the impact of three refining strategies on SeqRec under different iteration rounds, taking reflections on explicit user preferences as an example.
-From the Fig.
-3, we observed that:
+(a), (b), and (c) demonstrate the effects of iterations and different levels of refining on three datasets, while (d) shows the effects of the three perspectives with Games as an example.
+4.2. Recommendation Performance Comparison
+To answer RQ1, we compare the SeqRec performance of MoRE against baselines for 10 trials, with the average results presented in Tab.
+2 and Tab.
+3.
+We have the following observations: (1) MoRE outperforms all baselines across all metrics on all datasets, except for the suboptimal performance on the HR@1 on “Games”.
+(2) In general, LLM-based methods demonstrate performance comparable to the traditional methods, which validates the potential of LLMs in SeqRec.
+(3) Notably, MoRE outperforms LLM-based methods that specialize in predicting one single item (HR@1 only), even when those methods sacrifice ranking capabilities444 Note that these methods are inherently limited to top-1 predictions.
+Therefore, for these methods, we retain only top-1 results for a fair comparison here..
+In summary, MoRE exhibits superior performance in SeqRec, validating the efficacy of our approach.
+We attribute this to our proposed LLM-based reflection framework, which leverages LLM’s analytical and reasoning capabilities to capture users’ explicit preferences and implicit behavioral patterns while seamlessly integrating CF knowledge in a clear and comprehensible manner, ultimately enhancing performance in sequence recommendation tasks.
+Figure 5. Example of how LLM reflections from explicit preference (EP), implicit preference (IP), and collaborative filtering (CF) perspectives improve recommendation.
+4.3. Ablation Study: Multi-Perspective Reflection and Selection
+To validate the effectiveness of each perspective of reflection and answer RQ2, we conduct an ablation study on Amazon Arts. The results are presented in Tab. 4.
+We find that MoRE, equipped with the reflection perspective selection, outperforms alternatives, demonstrating the effectiveness of the multi-perspective reflections and the corresponding perspective selection.
+Specifically, MoRE outperforms the variants with not only the single reflection perspective but also the dual or triple perspectives simultaneously.
+The simultaneous use of multiple perspectives leads to a decrease in performance.
+This implies potential discrepancies among perspectives and no single perspective of reflection is universally optimal for all users, thereby confirming the necessity for selective choice among perspectives.
+Moreover, MoRE’s selection strategy surpasses both the Random and Greedy selection approaches which respectively denote a random reflection selection from the three memory banks and the selection solely based on downstream recommendation performance.
+This highlights the effectiveness of our contextual bandit modeling, enabling MoRE to choose the most suitable reflection for each user.
+4.4. Ablation Study: Refining and Iteration
+To answer RQ3, we test the impact of three refining strategies on SeqRec under different iteration rounds, taking reflections on explicit features as an example. From Fig. 4, we find:
 (1) Iteration: Overall, as the number of iterations increases, the recommendation performance gradually improves.
@@
 (2) Refining: Refining at group and individual levels shows a significantly better trend over that at the global level during iterations, particularly on Arts (Fig.
-3(a)).
-This improvement may be due to the global-level reflections introducing noise from other users, whereas group- and individual-level demonstrations more effectively elicit personalized reflective capabilities from the reflectors.
-Additionally, we found that the refining and iteration strategies are most effective for reflections on explicit preferences, compared to reflections on implicit preferences and collaborative filtering.
-Detailed experiments and discussions are presented in App.
-A.4.
-4.5. Analysis of LLM Training Cost
-To address RQ4, we conduct 5 repeat experiments to compare the training time and maximum memory consumption associated with all the LLM-based baselines that require training.
-To mitigate the effects of experimental randomness, we perform 5 repetitions of the experiments and report the results as the mean and standard deviation (SD), as shown in Tab.
-2.
-Our findings are as follows:
-(1) The training cost of reflection-based methods is lower than fine-tuning methods.
-This is because the reflection-based methods do not require fine-tuning, thus avoiding significant training time and memory costs.
-(2) Under reflection-based setup, MoRE has lower training time and memory consumption compared to Re2LLM.
-We attribute this to the larger action space of Re2LLM.
-Specifically, MoRE’s Meta-Reflector does not require selecting specific reflections but only needs to choose among three directions of reflection.
-In contrast, Re2LLM must first acquire a knowledge base of reflection and, once fixed, conduct PPO training of the retrieval agent within that knowledge base.
-This means it selects specific reflections from the entire knowledge base, with a larger action space.
-In summary, MoRE has the lowest training costs including time overhead and GPU memory usage among all the LLM-based baselines, demonstrating its training efficiency.
+4(a)).
+We assume that the global-level reflections may introduce noise from other users, whereas group- and individual-level demonstrations more effectively elicit personalized reflective capabilities from the reflectors.
+4.5. Analysis of Computational Cost
+To address RQ4, we compare training costs (time and GPU memory) across LLM-based baselines using five repeated experiments, reported as Avg.
+$\pm$ SD [^5] Average and standard deviation.
+in Tab.
+5.
+We find: 1) Reflection-based methods (e.g., MoRE) incur lower costs than fine-tuning approaches by avoiding LLM parameter updates; 2) MoRE outperforms Re2LLM in efficiency due to its smaller action space.
+MoRE’s Meta-Reflector selects from 3 reflection perspectives, while Re2LLM trains a PPO-based retrieval agent over a fixed base [^6] The size is 20., requiring larger action-space exploration.
+In summary, MoRE achieves the lowest training costs (time and memory) among all baselines, demonstrating superior training efficiency.
+4.6. Case Study
+We present a case study demonstrating reflections from EP, IP, and CF perspectives to enhance LLMREC (Fig.
+5).
+The three perspectives improve performance as follows: EP Reflection focuses on recent interaction trends (e.g., game genre preferences) to align with explicit preferences.
+This results in the target item being recommended at the 8-th position (e.g., Street Fighter 5), effectively capturing short-term user intentions.
+IP Reflection identifies patterns in historical attributes (e.g., multiplayer functionality, ”PlayStation” brand) to adapt to implicit preferences.
+This moves the target to the 6-th position by analyzing deeper user behavioral patterns.
+CF Reflection leverages collaborative filtering signals and emphasizes highly-rated items, reducing diversity while incorporating historical ratings and relevant sequels.
+This achieves the top-1 recommendation position.
+The Meta-reflector dynamically selects the optimal reflection (CF in this case), enabling MoRE to deliver optimal recommendations.
+The reflection mechanism bridges the semantic gap between LLMs and CF: LLMs utilize simple score comparisons instead of interpreting embeddings.
+This approach cost-effectively equips LLMs with domain-specific knowledge (e.g., CF signals) while mitigating their adaptation challenges in specialized domains.
 5. Conclusion
@@
 It selects the most appropriate perspective and corresponding reflections for each user’s current recommendation using a contextual bandit algorithm.
-Extensive experiments conducted on three benchmark datasets demonstrate that MoRE consistently outperforms traditional recommendation methods and LLM-based recommendation methods with minimal GPU memory usage and training time overhead.
-Appendix A Appendix: Experiment Details
-A.1. Datasets Statistics
-Arts Games Instruments
-# Users 55970 55145 27404
-# Items 22612 17287 10450
-Avg. len 8.80 9.01 8.41
-Sparsity 99.96% 99.95% 99.92%
-Table 4. Statistics of the 3 pre-processed datasets. Avg. len represents the average length of interacted items per user.
-The processed data statistics are summarized in Table 4.
-Note that “Arts”, “Games” and “Instruments” represent the abbreviations of “Arts, Crafts and Sewing”, “Video Games” and “Musical Instruments” respectively in the Amazon dataset (Amazon Review Data (2018)), where all data include reviews in the range May 1996 - Oct 2018 (Ni et al., 2019).
-A.2. Baselines Details
-We adopt three types of representative sequential recommendation methods as baselines to validate the effectiveness of our MoRE:
-Traditional deep learning recommendation models include:
-  • Caser (Tang and Wang, 2018) is a CNN-based method that models user-item interactions as sequences to capture both local and global patterns.
-  • GRU4Rec (Hidasi et al., 2015) is an RNN-based SeqRec model with GRU modeling the user behavior sequences.
-  • BERT4Rec (Sun et al., 2019) adopts a bidirectional Transformer model and incorporates a masked prediction approach to capture the patterns in item sequences.
-  • SASRec (Kang and McAuley, 2018) exploits a unidirectional Transformer-based neural network to model the item sequences and predict the next item.
-  • FDSA (Zhang et al., 2019) models both item-level and feature-level transition patterns through a dual self-attention mechanism.
-Training-free LLM-based methods, i.e., the prompt-based recommendation methods include:
-  • LLM4RS (Dai et al., 2023) investigates the potential of LLMs in enhancing recommendation capabilities by aligning it with traditional information retrieval (IR) ranking techniques, specifically point-wise, pair-wise, and list-wise ranking.
-  • LLMRank (Hou et al., 2024) leverages LLMs for zero-shot ranking in recommendation tasks, addressing challenges related to interaction order and bias through specialized prompting techniques.
-Training-required LLM-based methods, i.e., fine-tuning and reflection-based recommendation methods include:
-  • LC-Rec [^2] To ensure a fair comparison, we ensured that each model’s recommendation process for each user was based on the same item candidate set of size 50.
-However, since LC-Rec’s training is based on the full item set as the shared candidate set, it is challenging to guarantee that the model does not output items outside the size 50 candidate set for a user solely through the prompt.
-To address this, we implemented output constraints for the generation process of LC-Rec, ensuring that it outputs only the permitted 50 items for each user (Zheng et al., 2024) is a fine-tuning LLMREC method that incorporates collaborative filtering information by expanding the vocabulary space with special token embeddings.
-It is currently the state-of-the-art method for LLMREC in the full-candidate SeqRec setting.
-  • BinLLM [^3] BinLLM introduces collaborative filtering knowledge to LLMs from the perspective of input text.
-However, it can only predict the relationship between one user and one item at a time, necessitating multiple inferences to provide recommendations for a user.
-To ensure a fair comparison, we modified the format of its prompt and reproduced it to make it output a list.
-(Zhang et al., 2024) is a fine-tuning LLMREC method that incorporates collaborative filtering information by injecting user and item embeddings as binary text sequences into the LLM’s input, without modifying the model’s vocabulary space.
-It is currently the state-of-the-art method for LLMREC conditioned on the given user-item pair.
-  • Re2LLM (Wang et al., 2024) is a reflection-based LLMREC method.
-It first constructs a reflection knowledge base (KB) with domain knowledge, and then retrieves suitable text in the fixed KB to insert into the LLM’s input, adapting the LLM to session-based recommendation.
-(a) Explicit Preference (EP) Reflect
-(b) Implicit Preference (IP) Reflect
-(c) Collaborative Filtering (CF) Reflect
-Figure 4.
-Impact of iteration on reflections for explicit preferences, implicit preferences, and collaborative filtering information on recommendation performance.
-The experiments are conducted on the Amazon Games dataset, utilizing an individual-level refining strategy.
-A.3. Implementation Details
-Considering the cost and efficiency of LLMs, we perform random sampling from three datasets respectively.
-The principle for sampling users is to exclude those with excessively long interaction sequences before performing random sampling.
-For each dataset, we sample 1,000 users.
-Note that items are kept the same as the pre-processed datasets.
-[…]
-For our MoRE method, we train a DMF (Xue et al., 2017) model as the source of CF Reflection Expert’s CF score (Sec.
-3.3).
-The embedding size of this DMF model is 64 for both users and items.
-This DMF model is also used to implement the user clustering mentioned in the Refining process in Sec.
-3.4.
-DMF model provides embedding vectors for users in the three datasets, and these embeddings are used to cluster the users in each dataset.
-We define 20 clusters for each dataset, with the cluster ID being an integer in $[0,19]$, inclusive.
-For the Iteration process in Sec.
-3.4, we set the total iteration to 15, and impacts of Refining and Iteration can be found in Sec.
-4.4.
-A.4. Impact of Iteration on Different Perspective Reflections on the Amazon Games Dataset
-In Sec.
-4.4, using LLM reflections on explicit user preferences as an example, we conducted experiments that demonstrated the effectiveness of refining strategies in enhancing reflection quality and improving recommendations.
-In this section, we explore the extent to which the iteration strategy improves the quality of reflections on explicit preferences, implicit preferences, and collaborative filtering information.
-As can be seen from Fig.
-4, we have the following three observations.
-(1) In terms of the explicit and implicit user preferences, applying the iteration strategies improve the corresponding reflections.
-The results further verify the effectiveness of our reflection iteration strategy.
-(2) Applying iteration strategies to reflections on explicit preferences leads to greater improvements in recommendation performance compared to applying the same strategies to reflections on implicit preferences.
-We attribute these differences to the reasoning capabilities of the LLM.
-It is because understanding explicit user preferences is relatively easy for LLMs, as item titles directly reflect user preferences—such as when two products have highly relevant titles.
-On the other hand, implicit preferences are conveyed through product attributes, which poses greater challenges for LLMs to reason.
-(3) The improvement in the CF reflection through iteration strategies is not significant and shows noticeable fluctuations.
-We believe that even though we circumvent the issue of CF mathematical computations by introducing a tool model for the LLM, numerical comparisons still remain a limitation for the LLM.
-In short, the EP and IP reflections are more suitable for the refining and iteration than the CF reflections.
-A.5. Case Study
-In Sec.
-4.3, we conduct ablation studies to quantitatively verify the effectiveness of our reflection selection.
-In this section, we further validate it qualitatively through a case study.
-We illustrate the recommendation performances using reflections generated by MoRE and other strategies, shown in the green and brown boxes on the next page.
-The case in the brown box illustrates how the original LLMREC, without reflections, failed to hit the recommendation target.
-Meanwhile, among the various reflection selection methods, the Greedy strategy, which selects reflections based on their performance on the validation set, chooses the EP perspective, ultimately hitting the target at the 7-th position in the recommendation sequence.
-The Random strategy, on the other hand, randomly selects the IP perspective, which did not hit the recommendation target.
-As shown in the green box, MoRE, with the help of the selector, ultimately chooses the CF perspective, hitting the target at the 1-st position in the recommendation sequence.
-In this case, we observe that the user’s interaction history shows a clear preference for console and action games, with a specific but subtle interest in purchasing items related to console game brands.
-This preference was not correctly emphasized in the reflections generated from the EP and IP perspectives.
-Specifically, the IP reflection focuses too much on past preferences for the same brand (e.g., Nintendo games), overlooking potential shifts in the user’s future brand interests.
-In contrast, based on the Nintendo brand and action games, the CF reflection uses CF information to infer that the user might be interested in PlayStation action games.
-This approach accurately captured the user’s needs, leading to a correct recommendation.
-In short, in this case study, MoRE selects the most appropriate reflective perspective, which greatly assisted the recommendation process.
-Moving forward, we believe that employing a reflection mechanism can effectively bridge the semantic gap between LLMs and CF by allowing LLMs to leverage simple score comparisons rather than struggling with the understanding of embeddings.
-Invoking a straightforward model as a tool to obtain CF scores circumvents the challenge of LLMs understanding CF semantics, thereby endowing LLMs with CF knowledge.
-Moreover, utilizing reflection along with calling upon simple models offers a cost-effective solution to mitigate various issues LLMs face when adapting to and understanding specialized domains where they might otherwise struggle.
-Case Study: MoRE’s Reflection Selection Prompt: I’ve purchased the following products in the past in order: {Ben 10 Omniverse 2 - Nintendo Wii U;…}.
-And after last previous recommendation (maybe a failed trial), you give the following reflects:{$Ref_{i}$}.
-Now there are $|\mathcal{C}_{u}|$ candidate products that I can consider to purchase next:$\mathcal{C}_{u}$}.Please rank these $|\mathcal{C}_{u}|$ products by measuring the possibilities that I would like to purchase next most, according to the given purchasing records and reflects.
-Please think step by step.
-Please show me your top 10 ranking results with order numbers.
-Split your output with line break.
-You MUST rank the given candidates.
-You can not generate candidates that are not in the given candidate list.
-Answer by MoRE perspective selection, where selected reflection (CF reflection) $Ref_{i}=Ref_{\mathrm{CF}}$ = ‘PREFERENCE: The user tends to prefer gaming-related items with high CF scores, particularly those from PlayStation 4 and Xbox 360 platforms, as well as items with high CF scores from previous interactions.
-ANALYSIS: Although the actor’s previous recommendations are diverse, some items such as {HORI Fighting Commander 4 Controller for PlayStation 4/3, …}, was not included in the actor’s recommendations.
-This suggests that the actor may have overlooked strong CF signals in favor of more diverse recommendations.
-SUGGESTIONS: Prioritize gaming-related items with high CF scores, especially those from PlayStation 4 and Xbox 360 platforms, as well as items with high CF scores from previous interactions.
-Consider including items with CF scores above 0.5, such as {}, to improve the accuracy of the recommendations.’:
-1.
-Street Fighter V - PlayStation 4 Standard Edition (CF score: 0.83), 2.
-Mass Effect 2 - PC (CF score: 0.78), 3.
-Assassin’s Creed Liberation HD [Download] (CF score: 0.76), 4.
-Metal Gear Solid HD Collection (CF score: 0.75), 5.
-Call of Duty: Advanced Warfare Atlas Limited Edition - Xbox One (CF score: 0.74), 6.
-Spec Ops: The Line - Playstation 3 (CF score: 0.73), 7.
-Cities: Skylines [Online Game Code] (CF score: 0.72), 8.
-Final Fantasy Crystal Chronicles: Echoes of Time - Nintendo Wii (CF score: 0.71), 9.
-Fatal Fury: Mark of The Wolves (CF score: 0.7), 10.
-.hack, Part 3: Outbreak (CF score: 0.69)
-Case Study: Other Reflection Selectors Prompt: (same form as former prompt)
-Answer by recommendation w/o reflection, where selected reflection (EP reflection) $Ref_{i}$ = ‘’:
-Didn’t hit at any position) 1.
-Final Fantasy Legend, 2.
-Assassin’s Creed Liberation HD [Download], 3.
-Spinido iPhone 7 Plus Desk Stand Holder Compatible with All Smartphone with Cases- Gold, 4.
-Phineas and Ferb Ride Again - Nintendo DS, 5.
-Shadow Warrior, 6.
-Saints Row IV: National Treasure - Xbox 360, 7.
-Cabela’s African Adventures - Wii, 8.
-The Sims 2: Double Deluxe - PC, 9.
-Disney INFINITY (2.0 Edition) [Download], 10.
-Metal Gear Solid HD Collection
-Answer by Greedy perspective selector, where selected reflection (EP reflection) $Ref_{i}=Ref_{\mathrm{EP}}$ = ‘ PREFERENCE: The user has shown a strong interest in fighting games, particularly those with a classic or retro feel, as well as console-based games.
-ANALYSIS: The actor’s recommendations did not fully capitalize on the user’s preferences for fighting games and console-based games.
-The recommended games did not show a strong focus on gameplay and mechanics, and did not prioritize the most engaging and popular titles in the fighting game genre.
-SUGGESTIONS: Prioritize recommending fighting games, particularly those with a classic or retro feel, and consider suggesting console-based games that cater to the user’s interest in gameplay and mechanics.’:
-1.
-Fatal Fury: Mark of The Wolves, 2.
-Final Fantasy Crystal Chronicles: Echoes of Time - Nintendo Wii, 3.
-.hack, Part 3: Outbreak, 4.
-Capcom Classics Collection, 5.
-Bad News Baseball - Nintendo NES, 6.
-Teenage Mutant Ninja Turtles IV: Turtles in Time,
-7. Street Fighter V - PlayStation 4 Standard Edition, 8. Metal Gear Solid HD Collection, 9. Final Fantasy Legend, 10. Urban Champion
-Answer by Random perspective selector, where selected reflection (IP reflection) $Ref_{i}=Ref_{\mathrm{IP}}$ = ‘ PREFERENCE: Based on the user’s interaction history, it appears that they have a strong preference for games from Nintendo and PlayStation platforms, particularly games from the Wii and Wii U console platforms.
-The user also shows interest in games with multiplayer features, as evidenced by the inclusion of games like {}.
-ANALYSIS: To improve the actor’s recommendations, I suggest prioritizing games from Nintendo and PlayStation platforms, particularly games from the Wii and Wii U console platforms.
-The actor should also focus on games with multiplayer features, as this seems to be an area of interest for the user.
-Additionally, consider recommending games that offer unique gameplay mechanics, such as {}, which the user has interacted with previously.’:
-Didn’t hit at any position) 1.
-Final Fantasy Crystal Chronicles: Echoes of Time - Nintendo Wii, 2.
-Mike Tyson’s Punch-Out!!
-- Nintendo NES, 3.
-Bad News Baseball - Nintendo NES, 4.
-Urban Champion - Nintendo NES, 5.
-Super Mario 64 DS - Nintendo DS, 6.
-Moshi Monsters - Nintendo DS, 7.
-LEGO Pirates of the Caribbean - Nintendo DS, 8.
-Phineas and Ferb Ride Again - Nintendo DS, 9.
-Cabela’s African Adventures - Wii, 10.
-Disney Magical World - Nintendo 3DS
-Appendix B Appendix: Prompt design
-B.1. Base Recommendation Prompt
-We provide the base recommendation prompt as follows.
-Base recommendation prompt $P$ I’ve purchased the following products in the past in order:$\mathcal{S}_{u}$}.
-Now there are $|\mathcal{C}_{u}|$ candidate products that I can consider to purchase next: $\mathcal{C}_{u}$}.
-Please rank these $|\mathcal{C}_{u}|$ products by measuring the possibilities that I would like to purchase next most, according to the given purchasing records.
-Please think step by step.
-Please show me your top 10 ranking results with order numbers.
-Split your output with line break.
-You MUST rank the given candidates.
-You can not generate candidates that are not in the given candidate list.
-where $\mathcal{S}_{u}$ is user $u$’s previous interaction sequence in chronological order, $\mathcal{C}_{u}$ is the candidate item set for user $u$, and $|\mathcal{C}_{u}|$ is the candidate item set’s size.
-B.2. Recommendation with reflection’s prompt
-We provide the prompt of reflection-based recommendation as follows.
-Recommendation prompt with reflection $P_{\mathrm{REF}}$ I’ve purchased the following products in the past in order:$\mathcal{S}_{u}$}.
-And after last previous recommendation (maybe a failed trial), you give the following reflects:
-{$Ref_{i}$}.
-Now there are $|\mathcal{C}_{u}|$ candidate products that I can consider to purchase next:
-$\mathcal{C}_{u}$}.
-Please rank these $|\mathcal{C}_{u}|$ products by measuring the possibilities that I would like to purchase next most, according to the given purchasing records.
-Please think step by step.
-Please show me your top 10 ranking results with order numbers.
-Split your output with line break.
-You MUST rank the given candidates.
-You can not generate candidates that are not in the given candidate list.
-where $\mathcal{S}_{u}$ is user $u$’s previous interaction sequence in chronological order, $\mathcal{C}_{u}$ is the candidate item set for user $u$, and $|\mathcal{C}_{u}|$ is the candidate item set’s size.
-$Ref_{i}$ is the text of reflection from a specific perspective, including explicit preference (EP) reflection $Ref_{\mathrm{EP}}$, implicit preference (IP) reflection $Ref_{\mathrm{IP}}$ and collaborative filtering (CF) reflection $Ref_{\mathrm{CF}}$.
-B.3. Explicit Preference (EP) Reflector’s Generation Prompt
-In this section, we provide the detailed information about explicit preference (EP) reflector’s generation prompt, including zero-shot (EP reflection prompt $P_{\mathrm{EP}}$) and few-shot (EP reflection prompt $P_{\mathrm{EP}}$: With demonstrations) version.
-Here, we first show the zero-shot version, which is used for the beginning of the Refining and Iteration.
-EP reflection prompt $P_{\mathrm{EP}}$ You are an expert of giving reflections to the sequential recommender actor.
-The actor is a sequential recommender.
-It receives instructions and generates item recommendations based on user’s historical interactions.
-As an expert, you should provide reflections on the actor’s previous recommendations to help improve actor’s next recommendations for this user.
-You need to take it step by step.
-Step 1: Output 1 or 2 sentences starting with ’PREFERENCE:’.By analyzing the user’s interaction history based on text information such as item titles and categories with your knowledge.
-(paying more attention to later items in the history), you can predict the user preferences (e.g., trends, anomalies, interest).
-Step 2.
-Output 1 or 2 sentences starting with ’ANALYSIS:’.
-Then analyze the problems with the 10 items recommended by the actor based on user history and preferences.
-Step 3.
-Output 1 or 2 sentences starting with ’SUGGESTIONS:’.
-Finally provide suggestions for the actor to make next recommendations for this user (e.g.
-prioritize [specific aspect or feature]).
-Here are the actor’s recommendations you need to reflect on now (Don’t output any specific titles of items):
-Previous instruction:
-{$P$}
-Standard answer:$\hat{v}$} Actor’s recommendations:$\hat{\mathcal{O}}_{u}$}
-Please provide your response based on Previous recommendations below:
-where $P$ stands for the previous instruction, also known as the base recommendation prompt (see in App.
-B.1), which includes user’s previous interaction sequence and candidate item set.
-All items are displayed with their item titles, which represents the explicit preference of user.
-$\hat{v}$ is the standard answer of $P$, also known as the ground truth item which user will interacted with in the next time step.
-$\hat{\mathcal{O}}_{u}$ is actor’s recommendations, also known as the previous predicted item list including each item’s title.
-The task of $P_{\mathrm{EP}}$ is to generate the EP reflection $Ref_{\mathrm{EP}}$ based on all information mentioned above, especially the item titles which conveys the explicit preference.
-Then, we display the few-shot version, which is used for bettering Refining and Iteration’s performance by giving valuable demonstrations through specific strategies.
-EP reflection prompt $P_{\mathrm{EP}}$: With demonstrations You are an expert of giving reflections to the sequential recommender actor … Finally provide suggestions for the actor to make next recommendations for this user (e.g.
-prioritize [specific aspect or feature]).
-Here are some Examples:
-[Example 1 start]
-Previous instruction:
-{I’ve purchased the following products in the past in order: {Ancient Secrets: Quest for the Golden Key, … , Big City Adventure: Tokyo [Download]}.
-Now there are 50 candidate products that I can consider to purchase next: {SADES 7.1 Surround Sound Pro … White (SA903), The Jungle Book, … , Left 4 Dead - Game of the Year Edition -Xbox 360}.
-… Please show me your top 10 ranking results with order numbers.
-… }
-Standard answer:{Jewel Quest Mysteries: The Oracle of Ur [Download]} Actor’s recommendations:{[1.
-Dark Dimensions: Wax Beauty - PC, 2.
-Jewel Quest Mysteries: The Oracle of Ur [Download], … , 10.
-Dungeon Keeper 2 (Jewel Case) - PC]} Please provide your response based on Previous recommendations below:{”PREFERENCE: The user’s interest lies in games that involve puzzle-solving, adventure, and mystery, possibly with a focus on hidden object games, as they have interacted with similar games in the past.
-ANALYSIS: The actor’s recommendations did not fully capture the user’s preferences, instead suggesting a diverse mix of games across various platforms and genres.
-While the top 2 recommended items, Jewel Quest Mysteries: The Oracle of Ur [Download] and Dark Dimensions: Wax Beauty - PC, align with the user’s preferences, the subsequent recommendations stray away from the user’s interests.
-SUGGESTIONS: The actor should prioritize recommending games that fit the user’s preferences for puzzle-solving, adventure, and mystery, such as hidden object games or casual puzzle games, and focus on titles that have a similar tone and atmosphere to the user’s previously interacted items.} [Example 1 end]
-[Example 2 start]
-[Example 2 end]
-Here are the actor’s recommendations you need to reflect on … Please provide your response based on Previous recommendations below:
-B.4. Implicit Preference (IP) Reflector’s Generation Prompt
-In this section, we provide the detailed information about implicit preference (IP) reflector’s generation prompt, including zero-shot (IP reflection prompt $P_{\mathrm{IP}}$) and few-shot (IP reflection prompt $P_{\mathrm{IP}}$: With demonstrations) version.
-Here, we first show the zero-shot version, which is used for the beginning of the Refining and Iteration.
-IP reflection prompt $P_{\mathrm{IP}}$ You are an expert of giving reflections based on recommended items’ attributes to the sequential recommender actor.
-The actor is a sequential recommender.
-It receives instructions and generates item recommendations based on user’s historical interactions.
-As an expert, you should provide reflections based on the actor’s previous recommendations, especially items’ attributes to help improve actor’s next recommendations for this user.
-You need to take it step by step.
-Step 1: Output 1 or 2 sentences starting with ’PREFERENCE:’.
-By analyzing the user’s interaction history based on both text information such as items’ titles and items’ attributes with your knowledge.
-(paying more attention to LATER items’ attributes in the history), you can predict the user preferences (e.g., trends, anomalies, interest).
-Step 2.
-Output 1 or 2 sentences starting with ’ANALYSIS:’.
-Then analyze the problems with the 10 items recommended by the actor based on user history and preferences.
-Step 3.
-Output 1 or 2 sentences starting with ’SUGGESTIONS:’.
-Finally provide suggestions for the actor to make next recommendations for this user (e.g.
-prioritize [specific aspect, feature or attribute]).
-Here are the actor’s recommendations you need to reflect on now (Don’t output any specific titles of items):
-Previous instruction and items’ attributes (item titles, categories and brands):
-{$P$}
-Standard answer:$\hat{v}$} Actor’s recommendations:$\hat{\mathcal{O}}_{u}$}
-Please provide your response based on Previous recommendations below:
-where $P$ stands for the previous instruction, which includes user’s previous interaction sequence and candidate item set.
-Different from $P_{\mathrm{EP}}$, all items are displayed with their item titles and side information (including categories and brands), which represents the implicit preference of user.
-$\hat{v}$ is the standard answer of $P$, also known as the ground truth item which user will interacted with in the next time step.
-$\hat{\mathcal{O}}_{u}$ is actor’s recommendations, also known as the previous predicted item list including each item’s title.
-The task of $P_{\mathrm{IP}}$ is to generate the IP reflection $Ref_{\mathrm{IP}}$ based on all information mentioned above, especially the item’s side information which conveys the implicit preference.
-Then, we display the few-shot version, which is used for bettering Refining and Iteration’s performance by giving valuable demonstrations through specific strategies.
-IP reflection prompt $P_{\mathrm{IP}}$: With demonstrations You are an expert of giving reflections based on recommended items’ attributes to the sequential recommender actor.
-… Finally provide suggestions for the actor to make next recommendations for this user (e.g.
-prioritize [specific aspect or feature or attribute]).
-Here are some Examples:
-[Example 1 start]
-Previous instruction and items’ attributes (item titles, categories and brands):{I’ve purchased the following products in the past in order: {{title : “Red Heart Super Saver Yarn, Hunter Green”, categories : “Yarn”, brand : “Red Heart”}, … , {title : “Comfort Grip Crochet Hook 10/Pkg - Assorted Colors (one color per package)”, categories : “Crochet Hooks”, brand : “Kahoot”}}.
-Now there are 50 candidate products that I can consider to purchase next: {{title : “Liquitex Professional Gloss Gel, Medium, 8 Ounce (5708)”, categories : “Painting, Paints”, brand : “Liquitex”}} … Please show me your top 10 ranking results with order numbers.
-… }
-Standard answer:{Artgraf Watercolor Graphite In Tin}
-Actor’s recommendations:{[1. Sizzix 660097 Thinlits Die Magical Butterfly by Pete Hughes, … , 10. addi Knitting Needle Linos Straight 8 inch (20cm) Size US 06 (4.0mm)]}
-Please provide your response based on Previous recommendations below: {PREFERENCE: Based on the user’s interaction history, it appears that the user is interested in crafting supplies, particularly related to painting, drawing, and beading.
-ANALYSIS: The actor’s previous recommendations seem to be a mix of crafting supplies, but they lack a clear focus on the user’s preferences for painting, drawing, and beading.
-The recommendations also do not prioritize the user’s interaction history, which suggests a strong interest in crafting supplies.
-SUGGESTIONS: To improve the next recommendations, I suggest prioritizing items that are related to painting, drawing, and beading, as well as considering items that are similar to the user’s previous interactions.
-The actor should also consider recommending products from brands that are well-known for their quality and relevance to the user’s interests.}
-[Example 1 end]
-[Example 2 start]
-[Example 2 end]
-Here are the actor’s recommendations you need to reflect on … Please provide your response based on Previous recommendations below:
-Note that implicit preference (side information) are provided as the tuple “{title: …, categories : …, brand: …}” in user’s previous interacted sequence and candidate item set.
-B.5. Collaborative Filtering (CF) Reflector’s Generation Prompt
-In this section, we provide the detailed information about collaborative filtering (CF) reflector’s generation prompt, including zero-shot (CF reflection prompt $P_{\mathrm{CF}}$) and few-shot (CF reflection prompt $P_{\mathrm{CF}}$: With demonstrations) version.
-Here, we first show the zero-shot version, which is used for the beginning of the Refining and Iteration.
-CF reflection prompt $P_{\mathrm{CF}}$ You are an expert of giving reflections based on user & item’s collaborative filtering scores to the sequential recommender actor.
-The actor is a sequential recommender.
-It receives instructions and generates item recommendations based on user’s historical interactions.
-Then, a collaborative filtering model will provide scores based on current user and generated items.
-As an expert, you should provide reflections on the actor’s previous recommendations, especially user-item collaborative filtering scores to help improve actor’s next recommendations for this user.
-You need to take it step by step.
-Step 1: Output 1 or 2 sentences starting with ’PREFERENCE:’.
-By analyzing the user’s interaction history based on both text information such as item titles and categories, and historical items’ collaborative filtering scores with your knowledge.
-(paying more attention to items with HIGHER collaborative filtering scores in the history),you can predict the user preferences (e.g., trends, anomalies, interest).
-Step 2.
-Output 1 or 2 sentences starting with ’ANALYSIS:’.
-Then analyze the problems with the 0 items recommended by the actor based on user history and preferences.
-Step 3.
-Output 1 or 2 sentences starting with ’SUGGESTIONS:’.
-Finally provide suggestions for the actor to make next recommendations for this user (e.g.
-prioritize [specific aspect or feature]).
-Here are the actor’s recommendations you need to reflect on now (Don’t output any specific titles of items):
-Previous instruction and items’ collaborative filtering scores (item titles & CF scores):
-{$P$}
-Standard answer:$\hat{v}$} Actor’s recommendations:$\hat{\mathcal{O}}_{u}$}
-Please provide your response based on Previous recommendations below:
-where $P$ stands for the previous instruction, which includes user’s previous interaction sequence and candidate item set.
-Different from $P_{\mathrm{EP}}$ and $P_{\mathrm{IP}}$, all items are displayed with their item titles and collaborative filtering scores (the CF score, provided by a CF model), which represents the collaborative filtering information of user and items.
-$\hat{v}$ is the standard answer of $P$, also known as the ground truth item which user will interacted with in the next time step.
-$\hat{\mathcal{O}}_{u}$ is actor’s recommendations, also known as the previous predicted item list including each item’s title.
-The task of $P_{\mathrm{CF}}$ is to generate the CF reflection $Ref_{\mathrm{IP}}$ based on all information mentioned above, especially the user-item collaborative filtering information.
-CF reflection prompt $P_{\mathrm{CF}}$: With demonstrations You are an expert of giving reflections based on user & item’s collaborative filtering scores to the sequential recommender actor … Finally provide suggestions for the actor to make next recommendations for this user (e.g.
-prioritize [specific aspect or feature]).
-Here are some Examples:
-[Example 1 start]
-Previous instruction and items’ collaborative filtering scores (item titles & CF scores):{I’ve purchased the following products in the past in order: {{title : “Red Heart Super Saver Yarn, Earth and Sky”, CF score : “0.99847”}, … , {title : “Red Heart Super Saver Yarn, Light Sage”, CF score : “0.99846”}}.
-Now there are 50 candidate products that I can consider to purchase next: {{title : “Sizzix Bigz Die, 5.5 by 6-Inch, Picture Wheel”, CF score : “0.5”}, … , {title : “RED HEART E302C.0312 Super Saver Jumbo Yarn, Black”, CF score : “0.906811”}}.
-… Please show me your top 10 ranking results with order numbers.
-… }
-Standard answer:{RED HEART E302C.0312 Super Saver Jumbo Yarn, Black}
-Actor’s recommendations:{[1. Red Heart Super Saver Yarn, Earth and Sky, … , 10. Red Heart Super Saver Yarn, Melonberry Print]}
-Please provide your response based on Previous recommendations below: {PREFERENCE: Based on the user’s interaction history, it appears that they tend to prefer yarn and fabric related items with high CF scores.
-The user’s CF scores are highest for Red Heart yarn products, indicating a strong interest in these types of items.
-ANALYSIS: The actor’s previous recommendations did not fully leverage the user’s preferences, prioritizing items with lower CF scores.
-The recommended items, although diverse, did not seem to align with the user’s interests in crafts, DIY projects, and art supplies.
-SUGGESTIONS: To improve the next set of recommendations, I suggest prioritizing items that align with the user’s preferences and CF scores.
-The actor should consider recommending items related to crafts, DIY projects, and art supplies, such as yarn, fabric, and art materials.}
-[Example 1 end]
-[Example 2 start]
-[Example 2 end]
-Here are the actor’s recommendations you need to reflect on … Please provide your response based on Previous recommendations below:
-Note that collaborative filtering information are denoted as the tuple “{title: …, CF score: …}” in user’s previous interacted sequence and candidate item set.
+Furthermore, MORE’s unique self-improving meta-reflector, through its rigorous refining and iteration strategies, explicitly addresses the critical challenge of ensuring the validity and quality of LLM-generated reflections, providing reliable hints for recommendation.
+This mechanism effectively bridges the semantic understanding of LLMs with established recommendation principles, enhancing both performance and interpretability.
+Extensive experiments conducted on three benchmarks demonstrate that MoRE consistently outperforms traditional recommendation methods and LLM-based recommendation methods with minimal GPU memory usage and training time overhead.